A Bibliography of Publications in Fundamenta Informaticae: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.01

Title word cross-reference

(0, 1) [265]. 0L [221]. 1 [474]. 2 [105]. 3 [352, 156]. ϵ [434]. H [314, 350]. k [221]. λ [373, 111]. P [506]. p^{m} [494]. Y [307, 21]. ζ [355].

-Calculus [373, 111]. -Limited [221]. -logic [21]. -Optimal [434]. -Safe [474]. -Semantics [355]. -Space [156]. -Structures [105]. -Totality [265]. -Tree [307]. -valued [352].

0 [198].

49er [337].

Abductive [264]. Abilities [407]. Absolute [297]. Abstract [354, 176, 236, 533, 351]. Abstraction [364, 391]. Abstractness [545]. Acceptance [287, 543]. Accepting [504]. Access [484]. According [368].

Account [329]. ACI [198]. Acquisition [320]. Action [359, 318, 101, 545, 31]. Actions [192]. Activation [462]. Active [173]. Acyclic [163]. Adaptation [432]. Adapting [129]. Adaptive [420]. adders [30]. Adding [443]. Adequacy [170]. Adequate [402]. Adjacency [486]. Adjunctions [143]. Agent [524]. Agents [407, 50]. Ajdukiewicz [322]. Algebra [297, 508, 510, 478, 362, 359, 146, 164, 93]. Algebraic [281, 207, 300, 512, 15, 215, 274, 475, 396, 244, 275, 276, 518, 259, 8, 63, 91]. Algebras [487, 338, 505, 133, 362, 209, 230, 305, 254, 364, 391, 343, 388, 519, 67, 96, 13]. Algorithm [155, 437, 435, 463, 458, 92]. Algorithmic [201, 99, 26]. Algorithms [471, 432, 451, 461, 430, 309, 476, 434, 518, 334, 365, 294]. Along [372]. Alphabet [313, 56]. Alphabets [445]. AltaRica [537]. Ambiguities [323]. Ambivalent [268]. among [505]. Analyse [295]. Analysis [389, 251, 205, 425, 424, 458, 464, 427]. Analytical [258]. Analyzing [477]. Anomalies [162]. Answering [332]. Answers [368]. Ant [439]. Antichains [223]. Application [511, 235, 245, 458, 42]. Applications [508, 155, 157, 26, 293]. Applying [541]. Approach [508, 168, 437, 516, 205, 292, 351, 291, 337, 30, 50, 63, 91, 53]. Approaches [374]. Approximate [120, 165, 249, 332]. Approximating [273]. Approximation [257, 26]. Approximations [389, 370]. Arbitrary [181, 265]. Architecture [469]. Arcs [539, 542]. Argumentation [289]. Arithmetic [1, 40, 245]. Array [245]. Array-Arithmetic [245]. Artificial [466, 48]. Asking [311]. Aspects [439, 300, 75]. Assignment [279, 229]. Assistant [524]. Association [546, 382]. Assumptions [102, 371]. Asymptotic [109]. Asynchronous [208, 541, 239]. ATMS [461]. ATMS-based [461]. atomic [65]. Atomicity [308, 101]. Attribute [300, 110]. Attributed [240]. autoepistemic [71, 19, 55, 47, 76, 77]. Automata [449, 212, 208, 307, 497, 489, 490, 315, 445, 38, 9, 25]. Automated [516, 315, 388, 180, 261]. Automatic [320]. Automating [274, 517]. Average [455, 525, 348, 18]. Average-Case [455]. Axiom [32]. Axiomatic [451, 404]. Axiomatization [401, 226, 379]. Axiomatizations [189, 131, 204]. Axioms [107, 92].

Back [530]. Bag [186]. Base [283]. Based [168, 128, 326, 306, 172, 414, 341, 161, 359, 467, 238, 259, 335, 294, 544, 461, 420]. Basic [339, 309, 515, 14]. Basis [282, 468]. BDI [354]. Behavior [240]. Behaviour [290]. behaviours [17]. Belief [267, 406, 283, 372]. Beliefs [310, 50]. Between [447, 469, 176, 244, 143, 533, 278, 316]. Bi [374]. Bi-models [374]. Bilinear [141]. Binary [494, 521, 18]. Bisimulation [243, 226, 379, 488, 101, 60]. Block [360]. Blocking [216]. Boolean [67, 43, 152, 148]. Bottom [514]. Bottom-up [514]. Bound [234]. Bounds [211, 253, 348, 349]. Bracketed [501]. Branching [525, 226, 379]. Building [415, 315]. Bush [49]. calculate [68]. Calculi [308, 514, 322, 83]. Calculus [373, 263, 479, 480, 230, 111, 345, 452, 364, 391, 87]. Can [123]. Canonical [181, 338]. Can’t [123]. Capacity [149, 56]. Cartesian [202]. Case [455, 326, 310, 59]. Case-Based [326]. Categorical [402, 363, 145, 58]. Categories [184, 143, 255, 360, 361, 147]. Categorisation [262]. causal [79, 82]. Causality [500]. Cause [385, 384, 479, 357]. Cause-Effect [385, 384, 479, 357]. CCS [60, 107]. Cecylia [366]. Cellular [490, 25]. Certain [149, 521, 481]. Chain [433]. Chains [223]. chameleonic [21]. Change [406, 264]. Channel [479]. Chaotic [462]. Characteristics [446]. Characterization [449, 113, 9]. characterizations [87]. Children [104]. Chomsky [307, 314]. Circuit [490, 344]. Circumscribed [284]. Circumscribing [310]. Circumscription [121, 120, 165, 444]. Clark [39]. Clashes [173]. Class [394]. Classes [149, 527, 481, 485, 163, 109, 39, 58, 14]. Classical [141, 129]. Classification [326, 150, 151, 437, 328, 336, 1, 40, 22]. classifier [469]. Clausal [345]. Clause [183]. Clause-Diffusion [183]. Clauses [514]. Clifford [512]. Closed [43, 371]. closure [5]. Co [329]. Co-Operative [329]. Coded [160]. Codes [494, 160]. Coding [508, 421, 419, 454]. Coherent [480]. coincides [20]. Coinductive [401]. Collage [213]. Combinational [344]. Combinatorial [320]. Combinators [145]. Combining [326, 512, 107, 371]. Common [259]. Communicating [13, 288]. Communication [479]. Commutative [339]. Comonads [139]. Compact [140]. compactness [54]. Comparison [260]. comparisons [23]. Comparitive [205]. Compatible [158]. Compiler [305]. Complete [226]. completed [3]. Completenesess [316]. Completeness [131, 278, 55]. Completion [515, 39]. completions [21]. Complexity [112, 455, 525, 134, 520, 511, 521, 321, 152, 214, 369, 205, 349, 485, 18, 81, 76]. Composed [252]. Composition [538, 202]. Compositional [108, 278, 316, 7, 27]. Comprehensive [355, 522]. Compression [423, 420, 426, 382, 458, 418]. Computable [389, 262, 365]. Computation [442, 540, 306, 429, 157, 252, 346, 433, 534]. Computational [212, 321]. Computed [234]. Computer [508, 510, 362]. Computing [441, 466, 512, 157, 148, 460, 440]. Concatenable [293]. Concept [251, 464, 337]. Conception [248]. Concepts [337, 10]. Concerning [380, 221, 499, 64]. Concurrence [200]. Concurrency [308, 541, 101, 61, 98]. Concurrent [537, 304, 296, 359, 417, 291, 14, 23, 24]. Conditional [132, 188]. Conditionals [48]. Conditions [235]. Configurations [159]. Confluence [176, 544]. conformisms [324]. congruence [60]. Congruences [313]. Consequence [377, 386, 115]. Conservativity [83]. Consistency [304, 55]. Constrained [421]. Constraint [416, 415, 172, 414, 516, 417, 412]. Constraint-Based [172]. Constraints [318, 51]. Constructing [303]. Construction [208, 494, 228]. Constructive [511]. Constructor [194]. Context [207, 168, 169, 339, 295, 265, 323, 173, 166, 420, 33, 56, 81]. Context-Based [168, 420]. Context-Free [207, 339, 265, 33, 56, 81]. Contexts [171, 94]. Contextual [394, 321, 501, 398, 216, 446, 180].

Contingency [337]. Contributions [528]. Control [432, 308, 341, 467]. Controlled [314, 350, 14, 23]. Controller [467]. Controllers [331]. Convergence [424, 458]. Convex [373, 156]. Cooperating [447, 149]. Coordinatizing [364, 391]. Cores [487]. Correctness [305]. Correspondence [499]. Cost [484, 454]. COSY [536]. Coupled [295]. Coupled-Context-Free [295]. CR [387]. CR-Structures [387]. Credibility [406]. Critical [116]. Crossing [159]. CSP [301]. CTL* [32]. Cubic [196]. Curves [518]. Cut [141]. cycle [515].

D0L [498]. Dangers [320]. Data [244, 409, 533, 256, 291, 336]. Database [262, 222, 3, 29]. Databases [531, 532, 43, 91]. Dataflow [146]. Daubechies [426]. DAWN [471]. Deadline [540]. Deadlock [474]. Decidability [213, 312, 499, 76]. decidable [99]. Decision [525, 461, 380, 329, 342, 137, 148, 205, 253, 348, 349, 258, 334, 410, 544]. Decision-Making [329]. Declarative [197]. Decompositions [487]. Deduction [183, 224, 388]. Deductive [531, 532]. Default [117, 168, 353, 120, 165, 369, 161, 106, 118, 292, 534, 41, 75, 22]. Defaults [116, 123]. Defeasible [122]. Defect [492]. Definability [472]. Defined [396, 190, 92]. Defining [354, 240]. Definitional [302]. Definitions [530]. Degenerate [140]. Degree [283]. Deletion [95]. Delivery [435]. Deluge [330]. Denotational [270, 61, 59]. Density [544]. Deontic [122]. Dependence [209, 309, 245, 63, 64, 94]. Dependencies [269, 300, 188, 91]. Dependent [142]. Depth [490, 148, 253, 348]. Derivation [216]. derivational [81]. derivations [33]. Deriving [318]. Describing [537]. description [529]. descriptions [59, 65]. Designing [269]. Designs [344]. Detail [420]. Detection [351]. Determination [469]. Deterministic [155, 205]. DF0L [498]. d’Horizon [433]. Diagnosis [344]. diagram [99]. Dialogues [173]. Diamond [36]. Different [156]. Differential [467]. Diffusion [183]. Dimensional [217]. Directions [319]. Discovery [291, 261]. Discrete [297]. Disjunctive [271, 532, 381, 106]. Dissemilattices [343]. Distance [156]. Distinguishability [489]. Distributed [447, 471, 183, 476, 417, 333, 440, 82, 52]. Distribution [378]. Distributions [455]. Distributive [211, 184, 111, 147, 15]. Diversity [177]. Division [360]. DKBS [332]. DNA [441]. Do [123, 275]. Domain [374, 444, 16]. Domain-theoretic [374]. Domains [186, 143]. Double [324]. Double-pushout [324]. Down [110, 514]. downward [5]. Driven [540]. DTOL [380]. Duadic [494]. Duration [452, 98]. Dyck [295]. Dynamic [60, 290, 92]. Dynamics [462].

E-cycle [515]. easy [68]. Ecological [463]. Economic [435]. Editor [124, 175, 138]. Editors [412, 366, 325]. Effect [385, 384, 479, 357]. Effective [444]. Efficient [394]. Elementary [542, 36]. Elements [197]. Elimination [141]. Embedding [41]. Empirical [259, 261]. entrenchment [71]. Entries [243]. entropic [13]. Entropy [421, 454]. Entropy-Constrained [421].

Enumerability [241]. Enumeration [494]. Epistemic [71, 161]. Equality [401]. Equation [459]. Equational [232, 509, 194, 164, 83]. Equations [478, 197, 375, 465, 491, 403]. Equivalence [225, 498, 493, 485, 79]. Equivalences [312]. errors [68]. ESM [236]. Establish [304]. Estelle [190]. ETOL [210]. Evaluation [298, 190, 144, 110]. Evaluations [111]. Even [203]. event [79, 24]. evidence [12]. Evidential [244]. Evolution [438, 467, 436]. Evolutionary [432, 442, 431, 430, 429, 435, 463, 433, 440]. Evolving [305]. Examination [116]. Examples [328]. Exclusion [547]. Executable [438]. Executing [413]. existence [22]. Existential [109, 365, 89]. Exits [243]. Expanding [241]. expansions [72]. Experimental [291]. Explanation [335]. Explicit [106, 145, 61]. Exploitation [431]. Exploration [431]. Explosion [320]. Expressibility [135]. Expressible [491]. Expressing [125]. Expressions [225, 533]. Expressive [449]. Extended [286, 86, 93]. Extending [508]. Extensible [503]. Extension [250, 345, 516]. Extensions [487, 298, 118, 192, 534]. Extraction [409].

Factors [395]. Facts [330, 335]. failure [42]. Fairness [153, 53]. Families [195]. Family [188]. Filters [426]. Finding [128, 475]. Finite [509, 212, 217, 497, 433, 99, 16, 26]. Finitely [313]. First [135, 249, 292, 224]. First-Order [135, 249, 224]. Five [327]. Fix [147]. Fix-Point [147]. Fixed [478, 157]. Fixed-Points [157]. Fixpoint [224]. flowchart [8]. foldings [319]. Fork [362]. Form [302]. Formal [248, 251, 329, 214, 344, 174]. Formalising [407]. Formalism [537, 6]. Formalization [167, 19, 287, 244]. Forms [478]. Formulae [452]. Formulas [284, 521, 129, 99]. Fossil [429]. Foundation [337, 8]. Foundations [412, 294, 15]. Founded [353, 20]. Four [263, 505]. Fractal [424]. fragments [96]. Frames [452]. Framework [127, 215, 516, 259, 90, 21]. Free [207, 339, 295, 491, 265, 323, 33, 56, 81]. Full [545]. fully [93]. Function [423, 462]. Functional [355, 359]. Functions [132, 234, 156, 152, 148, 247]. Further [66]. Fuzzy [219, 465, 331, 467, 37].

Gains [527]. Game [411]. Gaussian [455]. General [540, 444, 227, 486]. Generalisation [201]. Generalised [538]. Generalization [210, 380]. Generalized [377, 225, 395, 322, 547, 38, 319, 11]. Generated [213, 313]. Generation [211, 152, 546]. Generative [149, 56]. Generic [121]. Genetic [437, 463, 434, 294]. Genomes [450]. Geometric [213, 155, 512, 157]. Given [408]. Global [356, 205]. Gorbachev [49]. Graded [377]. Grammar [447, 149, 195, 185, 288]. Grammars [213, 160, 339, 235, 322, 501, 265, 323, 397, 398, 237, 110, 216, 56, 81, 11]. Grammatical [160, 203, 81]. Graph [232, 455, 233, 235, 481, 237, 392, 424, 241, 239, 240]. Graphs [196, 208, 358, 290]. Ground [411, 459]. grounded [72]. Groups [493]. Growth [234]. GSP [187]. Guest [412, 325, 124, 366, 138]. Guided [128].

Hand [479]. Hand-Shake [479]. Hard [273]. Hausdorff [140]. headed [442]. Herbrand [272]. Hierarchies [123, 337]. Hierarchy [483, 307, 341, 314, 65, 33]. High [538]. High-level [538]. Higher [301, 57]. higher-order [57]. Historical [438]. History [79, 429, 178]. Horn [284]. Hybrid [509]. hypergraph [324]. Hyperplanes [410]. Hypothesis [286, 119]. Hypothesization [169].

Idea [471]. Identical [408]. Identifying [485]. ignorance [49]. II [402, 151, 73, 23]. ILF [471]. Image [421, 419, 326, 423, 426, 382, 458, 418]. Imagination [174]. Implicational [188]. Improving [326]. Incomplete [503, 332, 90]. Incompleteness [193]. inconsistent [21]. induced [32]. Induction [194, 340]. Inductive [530, 227, 517, 70]. Inequalities [277]. Inference [530, 120, 165, 203, 129, 85, 86, 81]. Infinitary [17]. Information [389, 300, 406, 285, 252, 255, 63, 52, 69, 90]. Information-Theoretic [389]. inheritance [19]. Inhibitor [539, 542]. Injectivity [495]. Input [113]. Instance [128]. Instance-Based [128]. Integer [497]. Integral [465]. Integrating [476]. Integration [510]. intelligence [48]. Intelligent [327]. Intensional [126]. Interleaving [312]. Interpret [304]. Intersection [229]. Interval [189, 464, 260]. Interval-Set [260]. Introduction [114]. introspection [74, 50]. Intuitionistic [302, 511]. Invariants [318]. irreducible [9]. ISMIS’93 [124]. Isomorphism [243, 84]. Issue [114, 460, 418]. Iterated [462]. Iteration [226, 481]. Iterative [522, 423].

Jumping [383].

Kernel [361]. Kleene [104]. Knapsacks [495]. Know [526]. Knowing [127]. Knowledge [169, 327, 320, 172, 283, 126, 330, 331, 496, 46]. Kripke [193].

Labelled [363]. Labelling [179]. Lambda [140, 87, 83, 364, 391]. Lambek [322]. Lamport [192]. Language [195, 172, 225, 298, 214, 190, 346, 454]. Languages [304, 207, 354, 472, 182, 394, 218, 210, 215, 295, 211, 321, 341, 396, 481, 541, 203, 370, 222, 35, 66]. Large [485]. Last [109]. lattice [15]. Lattices [373, 125]. Lauer [536]. Laws [261]. Learn [526]. Learning [68, 162, 437, 468, 342, 425, 259, 336]. left [9]. left-linear [9]. Lemmata [481]. Length [494]. LERS [340]. Less [526]. Letter [313, 56]. level [538]. Limited [221]. Limits [397]. Lindenmayer [447, 504]. Lindström [372]. Line [408]. Linear [139, 443, 197, 375, 155, 473, 187, 502, 396, 203, 301, 462, 9, 65, 51]. Lines [372]. Lingual [480]. Linguistic [283]. Link [305]. Lists [451]. Local [134, 356]. Logic [373, 303, 267, 268, 139, 281, 352, 443, 282, 353, 120, 165, 127, 150, 151, 461, 520, 302, 473, 513, 362, 287, 474, 274, 135, 381, 413, 141, 344, 126, 249, 369, 161, 173, 106, 144, 301, 276, 192, 69, 467, 238, 119, 191, 299, 78, 5, 62, 39, 71, 44, 19, 15, 47, 34, 10, 75, 4, 90, 77, 21, 42, 22, 32]. Logical [282, 230, 59, 91]. Logics [132, 117, 356, 104, 502, 189, 131, 129, 204, 224, 544, 45, 73, 41, 54, 46, 16, 26, 70, 529]. Long [159]. looking [92]. Loop [351]. lot [435]. Lotos [298]. Lower [234, 186, 211]. Łukasiewicz [352].

Machine [336]. Machines [484, 365]. Making [329]. Mallat [427]. Managing [172]. Manipulation [493]. Many [45, 73, 502]. Many-Valued [502, 45, 73]. Map [462]. Maps [488]. Marcus [321]. Markov [433]. Matching [522, 86]. Mathematical [121, 258]. Matrices [317]. matrix [56]. Matroid [259]. Maximal [223, 384]. McCarthy [230]. Means [305]. Mechanisms [430]. Mechanization [301]. Membership [527]. Memory [366, 436]. Meta [268, 531]. Meta-Queries [531]. Meta-variables [268]. Metalogical [377]. Metamathematics [171]. Method [509, 338, 283, 310]. Methodology [303, 183]. Methods [432, 326, 424, 334, 259]. Métivier [208]. Metric [545]. Metrics [299]. Minimal [384, 253, 109, 459]. Minimal-Maximal [384]. Minimization [28]. Mix [396]. mixed [79]. mixed-ordering [79]. MLP [469]. Modal [376, 356, 272, 513, 287, 521, 502, 119, 42, 529, 544, 62, 45, 73, 75, 70]. modalizing [529]. Model [509, 116, 402, 128, 326, 273, 285, 474, 532, 230, 463, 315, 222, 453, 486]. Model-Based [326]. Modeling [461]. Modelling [406, 24, 293]. Models [132, 39, 338, 520, 227, 140, 364, 391, 291, 260, 1, 40, 374, 26]. moderately [72]. Modified [458]. Modifying [126]. Modularity [278, 316]. Modulo [243, 198, 9]. Monadic [523]. Monotone [397]. Monotonic [267, 127, 502, 46]. Moore [489]. Morphism [361]. Morphisms [499]. Morphology [258]. Multi [480, 501]. Multi-Bracketed [501]. Multiple [243, 172, 129]. Multiple-Valued [129]. Multiresolution [421, 427]. Multisets [451]. Mutations [450]. Mutual [547].

Names [403]. Narrowing [181]. Natural [346, 224]. Needs [368]. Negation [106, 371, 62, 42]. Negative [235]. negotiation [288]. Nelson [254]. nested [93]. Net [472, 542, 36]. Net-Definability [472]. Nets [538, 539, 339, 473, 474, 500, 475, 220, 107, 477, 163, 383, 278, 316, 293, 317, 446, 7, 27]. Network [448, 469]. Networks [468, 462, 146, 108]. Neural [468, 462]. Next [131]. Noetherian [223]. Non [267, 159, 127, 271, 312, 495, 502, 345, 222, 46]. Non-Clausal [345]. Non-Crossing [159]. Non-Disjunctive [271]. Non-Injectivity [495]. Non-Interleaving [312]. Non-Monotonic [267, 127, 502, 46]. Non-Procedural [222]. Noncircular [110]. Nondeterministic [205, 24]. Nondivisibility [286]. noninterleaving [98]. Nonlinear [467]. Nonmonotonic [287, 115, 41, 42]. Nonsequential [306]. Normal [282, 478, 150, 151, 302, 369]. Note [208, 228, 203, 325, 255, 138, 109, 79, 35, 49, 10]. Notes [63]. Nothing [375]. Notion [199, 200]. Notions [377, 92]. null [93]. Numbers [152].

Obituary [130]. Object [473, 390, 240]. Objects [121, 503]. observations [82, 17]. Oja [458, 425]. Oja-RLS [458]. one [56]. one-letter [56]. Only [127, 547]. Open [516, 488]. Operation [396]. Operational [61, 59, 405, 240]. Operations [538, 450, 271]. Operative [329]. Operator [424]. Operators [264, 11]. Opportunities [407]. Optimal [139, 434, 410, 454, 378]. Optimization [137]. Optimizations [453]. OR-Graphs [290]. Order [135, 249, 358, 301, 292, 224, 109, 294, 57, 523]. Order-Based [294]. Ordering [263, 79]. Orderings [177, 178]. Orders [443]. ordinals [5]. other [50]. Output [113]. Overview [432].

Papers [124, 166]. Parallel [339, 242, 288, 440]. Parallelism [398, 216, 97]. Parameter [432]. Parameterized [405]. Parameters [184]. Parametric [297, 142]. Parametricity [404]. Parametrization [184]. parsable [35]. Parser [394]. parsimonious [72]. Parsing [172, 110]. Part [14, 23]. Partial [57, 443, 263, 358, 347, 67, 88]. Partitioned [423]. Path [543]. Paths [390, 477]. Pattern [323, 409, 86]. PCF [402]. Transition [539, 27]. Perfect [520, 1, 40, 74]. Performance [298, 190]. Periodic [540]. Perspective [438, 173, 276]. Peter [536]. Petri [538, 339, 473, 474, 7, 220, 107, 477, 163, 383, 293, 317]. PF [35]. Pictures [213]. Piecewise [462]. Place [539]. Place/Transition [539]. Plane [159]. Point [478, 327, 147]. Points [157]. poly [87]. poly-time [87]. Polyhedron [196]. polymorphic [88]. Polymorphism [142]. Polynomial [493, 365]. Polynomials [286, 464]. Pomsets [293]. Poset [277]. Positive [279]. Possibilistic [461, 44]. Possible [117]. Post [499]. Power [449, 212, 210, 380, 498, 491, 241, 278, 316]. Power-Free [491]. Powerdomain [373, 228, 390]. Practical [51, 403]. Practice [166]. Pragmatic [508]. Praise [268]. PRAM [456]. Precise [483, 93]. Precision [250]. predicate [27]. predicate/transition [27]. Prediction [326]. Preface [412, 470, 154, 428, 231, 400, 535, 482, 366, 393, 266, 280, 460, 175, 507, 100, 206, 418, 246, 124]. Preference [116]. Preferences [368]. Preferential [132]. Prefix [226]. Pregrammars [399, 457]. Preserving [420, 79]. Primitive [105, 89]. Principles [367]. Priorities [500]. Priority [153]. Probabilistic [169, 327, 237, 99]. Probabilities [109]. Problem [170, 455, 415, 527, 498, 435, 203, 542, 192, 499, 49, 64]. Problems [474, 380, 137, 351, 499, 85]. Procedural [222]. Procedures [544]. Process [297, 539, 263, 505, 472, 308, 312, 359, 467, 6]. Processes [243, 233, 339, 379, 187, 357, 358, 108, 293, 317, 446, 82, 24, 13]. Processing [532, 52, 37]. Processors [408]. Procrastination [530]. Product [360]. production [37]. Productions [397]. Products [145]. Program [511, 245, 125]. Programmed [238]. Programming [267, 268, 416, 304, 197, 414, 276, 417, 412, 62, 15, 90, 78]. Programs [303, 282, 525, 355, 150, 151, 520, 437, 474, 135, 381, 413, 347, 106, 191, 299, 543, 147, 1, 40, 5, 39, 92, 68, 15, 4, 21]. progressing [60]. Proof [416, 112, 511, 305, 28]. Proofs [286, 274, 517, 108, 278, 316, 164]. Propagation [376]. Properties [385, 201, 150, 151, 213, 327, 176, 523, 383, 36, 26]. Property [275]. Proposal [372]. Propositional [310, 344, 345, 192, 18, 54]. prove [42]. Provers [508, 510]. Proving [128, 512]. Prune [157]. Prune-And-Search [157]. Psychology [102]. Pure [504, 323, 397]. pushout [324].

Qualification [170]. Qualifying [368]. Qualitative [461, 80, 519]. Quantification [47]. Quantified [513]. quantifying [74]. quantifying-in [74]. Quantisation [421, 420, 422]. Quantitative [199]. Quantization [468]. Quasi [299]. Quasi-Metrics [299]. Queries [531, 162, 262, 365]. Query [52, 332]. Questions [311].

Rabinowicz [372]. Radial [468]. Radius [350]. Ramified [521]. Random [484]. Randomized [456]. Rank [229]. Rational [102, 187, 152, 396, 518]. Rauszer [366]. RCC [30]. RCC-adders [30]. Reachability [474, 163]. Read [539]. Real [298, 174, 301, 518]. Real-Time [298, 174, 301]. Reals [389]. Reason [403]. Reasoned [102]. Reasoning [267, 168, 170, 127, 72, 283, 502, 49, 249, 122, 292, 256, 335, 260, 180, 519, 19, 51, 50, 76, 21]. Recognizability [217]. reconstruction [88]. Recurrent [493]. recursion [34, 89]. Recursive [401, 227, 369, 241, 279]. reduced [66]. Reducing [271, 320, 11]. Reduction [284, 399, 336]. Reductions [139, 444, 105]. Reducts [252]. Refinement [125, 101, 545, 31, 65]. Reflective [365]. Reflexive [77]. Reformulating [468]. Regression [464]. Regular [379, 182, 225, 533, 499]. Rejection [287]. Relabeling [241]. Relate [132]. Relating [82]. Relation [469, 201, 133, 533, 382, 115, 96]. Relational [374, 361, 164, 43, 93, 91]. Relations [338, 356, 519, 176, 247, 403]. Relationship [447, 244]. Relationships [505, 278, 316]. Relative [297, 484]. Relevant [451]. Reliability [455]. Remarks [447, 450, 3, 66]. Representable [136]. Representation [338, 309, 358, 331, 317, 67]. Representations [327]. Representing [479, 283, 390]. Requirements [125]. Residuated [338]. Resolution [16, 28]. respect [99]. Response [540]. Restricted [176, 42]. Result [284, 193]. Results [304, 313, 433, 221, 178, 85]. Retrieval [92, 285]. Reversible [490]. Review [419, 441]. Revision [267, 372]. Revisited [385]. Revisiting [436]. Rewrite [181, 415, 182, 176]. Rewriting [232, 512, 501, 392, 238, 179, 180, 229, 9, 324]. Riemann [286]. RLS [425, 458]. roles [529]. Rough [281, 352, 248, 133, 285, 250, 251, 249, 528, 331, 546, 254, 370, 247, 334, 256, 453, 291, 464, 259, 388, 260, 10, 12]. Rough-Set [256, 260]. Rule [340, 414, 345, 546, 425, 42]. Rule-Based [414]. Rules [376, 415, 337].

Safe [474]. Satisfaction [415]. Satisfiability [277, 192]. SCCS [363]. Scheduling [540, 435]. scheme [8]. Schemes [306]. Science [362]. Scott [386]. Search [509, 157]. Second [523, 109]. Second-Order [109, 523]. Seeking [330]. Selected [124]. Selection [132, 430]. Self [432]. Self-Adaptation [432]. Semantic [448, 385, 245, 179, 31, 53]. Semantically [128]. Semantics [448, 304, 117, 353, 539, 355, 263, 296, 150, 151, 271, 306, 405, 193, 287, 273, 474, 135, 298, 500, 236, 532, 363, 106, 144, 359, 276, 191, 299, 543, 147, 101, 545, 62, 61, 2, 58, 7, 20, 27]. Semaphores [547]. Semi [284, 295]. Semi-Dyck [295]. Semi-Horn [284]. Semigroups [338]. Seminormal [534]. Semiorders [223]. Sentences [283, 245]. Separators [155]. Sequences [493]. Sequent [480]. Sequential [135, 329, 502]. serializability [29]. Series [136, 210, 380, 498]. Set [285, 370, 334, 256, 453, 291, 260, 99, 68]. Sets [281, 352, 248, 201, 451, 295, 250, 251, 341, 502, 249, 528, 546, 254, 247, 113, 464, 259, 499, 9, 43, 95, 12]. Setting [522, 329]. Shake [479]. shape [25]. Shift [271]. Shift-Operations [271]. Shifts [173]. Short [286]. Signal [290]. Signed [381, 129]. Silent [449]. Similarity [356, 185, 464]. Simple [402, 459, 82]. Simplification [515, 178]. Simulation [456, 489]. Simultaneous [110]. Size [445]. Skew [527]. Skolemization [112]. Slender [341]. Small [350]. Smyth [228]. Snack [343]. Soft [466, 460]. SoleX [516]. Solitaire [496]. Solution [170, 375]. Solve [474]. Solver [516]. Solvers [415]. Solving [493]. Some [439, 338, 377, 367, 320, 309, 3, 313, 253, 192, 351, 221, 163, 499, 39, 85]. Soundness [131]. Sources [172]. Space [156]. Spaces [309, 480, 502, 257, 64]. Sparse [286]. Spatial [290, 519]. Special [114, 460, 418]. Specific [383]. Specification [510, 298, 184, 190, 517, 6]. Specifications [523, 275, 240, 57]. Spoken [172]. Square [103]. Stability [425]. Stable [273, 474, 532, 502, 371, 79, 20]. Stack [445]. Stage [292]. Standards [419]. State [217, 406]. States [161]. Static [305]. Stationary [118]. Statistical [326]. Step [109]. Stochastic [458, 454]. Stopping [434]. Strategies [181, 415]. Stratification [134]. Stratified [2, 1, 40, 4]. Stream [202]. Strings [177]. Strong [150]. Strongly [223, 16]. Structural [376, 385, 347, 107]. Structurally [490]. Structure [214, 238, 99]. Structured [405]. Structures [438, 385, 384, 479, 105, 244, 342, 390, 387, 79, 24, 22]. Study [201, 423, 59]. Studying [388]. Subdirect [487]. Subgoal [514]. Subgroups [219]. Submonoids [457]. Subset [521]. Substitution [84, 242]. Substitutivity [367]. subterm [96]. subtypes [86]. Subtyping [503, 401]. Subwords [491]. Suggested [450]. Survey [334]. Suspended [413]. Switching [527]. Symbolic [197]. Symmetries [475]. Symposium [124]. Synchronization [218, 408, 220, 506, 445]. Synchronous [541, 239]. Syntactic [197, 245, 50]. Syntax [268, 298, 97]. Synthesis [511, 251, 347, 542]. System [375, 340, 475, 332, 404, 337, 92, 65, 41, 32]. Systems [447, 537, 510, 181, 504, 439, 149, 296, 386, 327, 300, 103, 187, 182, 176, 380, 236, 476, 252, 357, 190, 423, 185, 359, 488, 314, 350, 506, 204, 542, 255, 289, 333, 238, 241, 459, 335, 229, 9, 36, 14, 23, 63, 288, 52, 69, 221, 28, 37]. Systolic [307].

Tableaux [376, 544]. Tables [244, 258, 291, 337]. Tasks [540]. Teams [162]. Technique [305]. Techniques [320, 541, 129, 388]. Temporal [443, 34, 131, 204, 192, 82, 54]. Tennenbaum [275]. Tentative [157]. Term [232, 512, 459, 179, 229, 9, 96]. Term-Rewriting [512]. Terminally [160]. Termination [176, 392, 179]. Terms [201, 522, 43]. test [30]. Tetrahedralizations [158]. Text [215]. Their [104, 523, 293, 511, 446]. Theorem [483, 510, 128, 512, 272, 54]. Theorems [492]. Theoretic [416, 389, 334, 374]. Theoretical [294]. Theories [509, 367, 534, 318, 83, 55]. Theory [508, 207, 170, 150, 151, 214, 528, 166, 119, 372, 259, 80, 8, 12, 29]. Three [104, 2, 20]. three-valued [2, 20]. Thue [523, 313]. Tilson [360]. Time [297, 483, 525, 384, 540, 155, 298, 408, 349, 98, 434, 174, 301, 477, 365, 87]. Timed [449, 505, 296, 44, 379, 452, 488, 7]. Token [378]. Tolerance [257]. Tool [471]. tools [51]. Top [514, 110]. Top-Down [110, 514]. Topology [191]. Topos [363]. total [324]. Totality [265]. Tour [433]. Trace [481, 204]. Transactional [333]. Transducer [411]. Transducers [234]. Transformations [302, 511, 239, 240]. Transformer [202]. Transforming [358]. Transition [378, 488]. Transition-optimal [378]. Transitions [449]. Traps [487]. Tree [136, 234, 306, 307, 182, 489, 205, 253, 315, 445, 411, 9]. Tree-Based [306]. Trees [525, 160, 521, 492, 363, 137, 148, 348, 349, 410, 319]. Trends [440]. Triangulated [208]. Truth [502, 452]. TSP [439]. Two [217, 211, 313, 292]. Two-Dimensional [217]. Two-Letter [313]. Two-Stage [292]. Type [401, 85, 86, 142, 279, 229, 58, 88]. Type-Dependent [142]. typed [83]. Types [227, 533, 89].

Un- [213]. Unambiguous [533]. Unary [487]. Uncertain [256, 260]. Uncertainty [367, 328, 80]. Undecidability [497, 225, 313, 221, 88]. Undecidable [265, 96, 392]. Unearthing [429]. Unification [375, 451, 522, 198]. Unified [127, 29]. Uniform [451, 30]. Uniformity [199]. Unimprovable [349]. uniqueness [25]. Universal [222]. Universality [489]. unmarked [27]. Unseen [328]. Upper [349]. Use [514]. Useful [327]. User [368]. Using [169, 415, 461, 474, 295, 184, 129, 125, 324].

V [219]. valid [99]. Valuation [335]. Valued [104, 502, 129, 352, 2, 45, 73, 20]. Variable [250]. variables [268]. Varying [132]. Vector [421, 468, 14, 23, 422, 543]. Verifiability [311]. Verification [344, 245]. Verifying [471]. Version [483, 340]. Versus [374, 457, 311]. VGA [469]. VGA-classifier [469]. Via [162, 249, 107, 65]. Video [419]. View [416, 451, 327, 473]. Viewpoints [167]. Virus [463]. Virus-Evolutionary [463]. Vision [466]. vs [60].

Wavelet [420, 426]. Wavelet-based [420]. Wavelets [427]. Ways [403]. weak [2, 151]. Weakly [4]. Weighted [212, 497, 348, 293]. Well [353, 20]. Well-Founded [353, 20]. we’re [442]. Where [442]. Wille [94]. Without [131]. Word [491, 481, 242]. Words [395, 491]. Workflows [333]. World [371]. Worlds [117]. Wulff [25].

Zero [496]. Zero-knowledge [496].

References

Apt:1990:ACP

[1] K. R. Apt and H. A. Blair. Arithmetic classification of perfect models of stratified programs. Fundamenta Informaticae, 13(1):1–17, March 1990. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fitting:1990:SWS

[2] M. Fitting and M. Ben-Jacob. Stratified, weak stratified and three-valued semantics. Fundamenta Informaticae, 13(1):19–33, March 1990. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kunen:1990:SRC

[3] K. Kunen. Some remarks on the completed database. Fundamenta Informaticae, 13(1):35–49, March 1990. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Przymusinska:1990:WSL

[4] H. Przymusinska and T. C. Przymunsinski. Weakly stratified logic programs. Fundamenta Informaticae, 13(1):51–65, March 1990. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bagai:1990:DCO

[5] R. Bagai, M. Bezem, and M. H. van Emden. On downward closure ordinals of logic programs. Fundamenta Informaticae, 13(1):67–83, March 1990. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mauw:1990:PSF

[6] S. Mauw and G. J. Veltink. A process specification formalism. Fundamenta Informaticae, 13(2):85–139, June 1990. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maggiolo-Schettini:1990:CST

[7] A. Maggiolo-Schettini and J. Winkowski. A compositional semantics for timed Petri nets. Fundamenta Informaticae, 13(2):141–170, June 1990. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cazanescu:1990:TNA

[8] V. E. Căzănescu and Gheorghe Stefănescu. Towards a new algebraic foundation of flowchart scheme theory. Fundamenta Informaticae, 13(2):171– 210, June 1990. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fulop:1990:CIS

[9] Z. Fülop and S. Vágvölgyi. A characterization of irreducible sets modulo left-linear term rewriting systems by tree automata. Fundamenta Informaticae, 13(2):211–226, June 1990. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Krynicki:1990:NRC

[10] M. Krynicki. A note on rough concepts logic. Fundamenta Informaticae, 13(2):227–235, June 1990. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Novotny:1990:ROG

[11] M. Novotný. Reducing operators for generalized grammars. Fundamenta Informaticae, 13(3):237–244, September 1990. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skowron:1990:RST

[12] A. Skowron. The rough sets theory and evidence theory. Fundamenta Informaticae, 13(3):245–262, September 1990. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Romanowska:1990:CPE

[13] A. B. Romanowska and J. D. H. Smith. Communicating processes and entropic algebras. Fundamenta Informaticae, 13(3):263–274, September 1990. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Keesmaat:1990:VCC

[14] N. W. Keesmaat, H. C. M. Kleijn, and G. Rozenberg. Vector controlled concurrent systems. Part I: basic classes. Fundamenta Informaticae, 13(3): 275–316, September 1990. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hirani:1990:AFL

[15] A. Hirani and V. S. Subrahmanian. Algebraic foundations of logic programming, I: the distributive lattice of logic programs. Fundamenta Informaticae, 13(3):317–332, September 1990. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stachniak:1990:RDS

[16] Z. Stachniak and P. O’Hearn. Resolution in the domain of strongly finite logics. Fundamenta Informaticae, 13(3):333–351, September 1990. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Darondeau:1990:IBI

[17] P. Darondeau and B. Gamatie. Infinitary behaviours and infinitary observations. Fundamenta Informaticae, 13(3):353–386, September 1990. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Buning:1990:TAC

[18] H. Kleine Büning and U. Löwen. Towards average complexity of propositional binary. Fundamenta Informaticae, 13(3):387–399, September 1990. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gelfond:1990:FIR

[19] Michael Gelfond and Halina Przymusinska. Formalization of inheritance reasoning in autoepistemic logic. Fundamenta Informaticae, 13(4):403– 443, December 1990. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Przymusinski:1990:WFS

[20] Teodor Przymusinski. Well-founded semantics coincides with three-valued stable semantics. Fundamenta Informaticae, 13(4):445–463, December 1990. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Subrahmanian:1990:LFR

[21] V. S. Subrahmanian. Y -logic: a framework for reasoning about chameleonic programs with inconsistent completions. Fundamenta Informaticae, 13(4):465–483, December 1990. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhang:1990:CES

[22] Aidong Zhang and Wiktor Marek. On the classification and existence of structures in default logic. Fundamenta Informaticae, 13(4):485–499, December 1990. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Keesmaat:1991:VCC

[23] N. W. Keesmaat, H. C. M. Kleijn, and G. Rozenberg. Vector controlled concurrent systems, Part II: comparisons. Fundamenta Informaticae, 14 (1):1–38, January 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Loogen:1991:MNC

[24] Rita Loogen and Ursula Goltz. Modelling nondeterministic concurrent processes with event structures. Fundamenta Informaticae, 14(1):39–74, January 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wlaz:1991:UWS

[25] Pawel Wlaź. On uniqueness of the Wulff shape for cellular automata. Fundamenta Informaticae, 14(1):75–90, January 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stepaniuk:1991:AFM

[26] Jaroslaw Stepaniuk. Applications of finite models properties in approximation and algorithmic logics. Fundamenta Informaticae, 14(1):91–108, January 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Schettini:1991:CSU

[27] Andrea Maggiolo Schettini, G. Michele Pinna, and Józef Winkowski. A compositional semantics for unmarked predicate/transition nets. Fundamenta Informaticae, 14(1):109–128, January 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stachniak:1991:MRP

[28] Zbigniew Stachniak. Minimization of resolution proof systems. Fundamenta Informaticae, 14(1):129–146, January 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vidyasankar:1991:UTD

[29] K. Vidyasankar. Unified theory of database serializability. Fundamenta Informaticae, 14(2):147–183, February 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Becker:1991:UTA

[30] Bernd Becker and Uwe Sparmann. A uniform test approach for RCCadders. Fundamenta Informaticae, 14(2):185–219, February 1991. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Darondeau:1991:ASA

[31] Philippe Darondeau and Pierpaolo Degano. About semantic action refinement. Fundamenta Informaticae, 14(2):221–234, February 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Koutny:1991:ASI

[32] Maciej Koutny. Axiom system induced by CTL* Logic. Fundamenta Informaticae, 14(2):235–253, February 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Makinen:1991:HCF

[33] Erkki Mäkinen. A hierarchy of context-free derivations. Fundamenta Informaticae, 14(2):255–259, February 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kroger:1991:TLR

[34] Fred Kröger and Stephan Merz. Temporal logic and recursion. Fundamenta Informaticae, 14(2):261–281, February 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Balanescu:1991:NPK

[35] Tudor Balanescu and Marian Gheorghe. A note on PF(k) — parsable languages. Fundamenta Informaticae, 14(3):283–286, March 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hoodgeboom:1991:DPE

[36] Hendrik Jan Hoodgeboom and Grzegorz Rozenberg. Diamond properties of elementary net systems. Fundamenta Informaticae, 14(3):287–300, March 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stoeva:1991:PFP

[37] Stefka P. Stoeva. On processing of fuzzy production systems. Fundamenta Informaticae, 14(3):301–312, March 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barua:1991:GA

[38] Rana Barua. On generalized automata. Fundamenta Informaticae, 14 (3):313–322, March 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Baratella:1991:MCC

[39] Stefano Baratella. Models of Clark’s completion for some classes of logic programs. Fundamenta Informaticae, 14(3):323–337, March 1991. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Apt:1991:ACP

[40] Krzysztof R. Apt and Howard A. Blair. Arithmetic classification of perfect models of stratified programs. Fundamenta Informaticae, 14(3):339–343, March 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kaminski:1991:EDS

[41] Michael Kaminski. Embedding a default system into nonmonotonic logics. Fundamenta Informaticae, 14(3):345–353, March 1991. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Truszczynski:1991:MNL

[42] Miroslaw Truszczynski. Modal nonmonotonic logic with restricted application of the negation as failure to prove rule. Fundamenta Informaticae, 14(3):355–366, March 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jankowski:1991:CSB

[43] Andrzej Jankowski and Zbigniew Michalewicz. Closed sets of Boolean terms in relational databases. Fundamenta Informaticae, 14(3):367–385, March 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dubois:1991:TPL

[44] Didier Dubois, Jérǒme Lang, and Henri Prade. Timed possibilistic logic. Fundamenta Informaticae, 15(3–4):211–234, November/December 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fitting:1991:MVM

[45] Melvin C. Fitting. Many-valued modal logics. Fundamenta Informaticae, 15(3–4):235–254, November/December 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Parikh:1991:MNM

[46] Rohit Parikh. Monotonic and non-monotonic logics of knowledge. Fundamenta Informaticae, 15(3–4):255–274, November/December 1991. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Konolige:1991:QAL

[47] Kurt Konolige. Quantification in autoepistemic logic. Fundamenta Informaticae, 15(3–4):275–300, November/December 1991. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Horty:1991:CAI

[48] John F. Horty and Richmond H. Thomason. Conditionals and artificial intelligence. Fundamenta Informaticae, 15(3–4):301–323, November/ December 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kraus:1991:RAI

[49] Sarit Kraus, Donald Perlis, and John Horty. Reasoning about ignorance: a note on the Bush–Gorbachev problem. Fundamenta Informaticae, 15(3– 4):325–332, November/December 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maida:1991:SAI

[50] Anthony S. Maida, Jacques Wainer, and Sehyeong Cho. A syntactic approach to introspection and reasoning about the beliefs of other agents. Fundamenta Informaticae, 15(3–4):333–356, November/December 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Huynh:1991:PTR

[51] Tien Huynh, Leo Joskowicz, Catherine Lassez, and Jean-Louis Lassez. Practical tools for reasoning about linear constraints. Fundamenta Informaticae, 15(3–4):357–380, November/December 1991. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ras:1991:QPD

[52] Zbigniew W. Ras. Query processing in distributed information systems. Fundamenta Informaticae, 15(3–4):381–397, November/December 1991. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rutten:1992:SAF

[53] J. J. M. M. Rutten and J. I. Zucker. A semantic approach to fairness. Fundamenta Informaticae, 16(1):1–38, January 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Leischner:1992:CTP

[54] Martin Leischner. On the compactness theorem of propositional temporal logics. Fundamenta Informaticae, 16(1):39–57, January 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gelfond:1992:CCA

[55] Michael Gelfond and Halina Przymusinska. On consistency and completeness of autoepistemic theories. Fundamenta Informaticae, 16(1):59–92, January 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Makinen:1992:GCC

[56] Erkki Mäkinen. On the generative capacity of context-free matrix grammars over one-letter alphabet. Fundamenta Informaticae, 16(1):93–97, January 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Astesiano:1992:PHO

[57] Egidio Astesiano and Maura Cerioli. Partial higher-order specifications. Fundamenta Informaticae, 16(2):101–126, February 1992. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hilken:1992:TCS

[58] Barney P. Hilken and David E. Rydeheard. Towards a categorical semantics of type classes. Fundamenta Informaticae, 16(2):127–147, February 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Egidi:1992:ODL

[59] Lavinia Egidi, Furio Honsell, and Simona Ronchi Della Rocca. Operational, denotational and logical descriptions: a case study. Fundamenta Informaticae, 16(2):149–169, February 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Montanari:1992:DCV

[60] Ugo Montanari and Vladimiro Sassone. Dynamic congruence vs. progressing bisimulation for CCS. Fundamenta Informaticae, 16(2):171–199, February 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Broy:1992:ODS

[61] Manfred Broy. Operational and denotational semantics with explicit concurrency. Fundamenta Informaticae, 16(3–4):201–229, March 1992. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Balbiani:1992:MSN

[62] Philippe Balbiani. A modal semantics of negation in logic programming. Fundamenta Informaticae, 16(3–4):231–262, March 1992. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Novotny:1992:NAA

[63] Jiří Novotný and Miroslav Novotný. Notes on the algebraic approach to dependence in information systems. Fundamenta Informaticae, 16(3–4): 263–273, March 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Novotny:1992:PCD

[64] Miroslav Novotný and Zdzisław Pawlak. On a problem concerning dependence spaces. Fundamenta Informaticae, 16(3–4):275–287, March 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gorrieri:1992:HSD

[65] Roberto Gorrieri. A hierarchy of system descriptions via atomic linear refinement. Fundamenta Informaticae, 16(3–4):289–336, March 1992. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paun:1992:FRR

[66] Gheorghe Paun, Miklos Szijarto, and Sorina Vicolov. Further remarks on reduced languages. Fundamenta Informaticae, 16(3–4):337–347, March 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ehrenfeucht:1992:RPB

[67] Andrzej Ehrenfeucht and Marek W. Zawadowski. A representation of partial Boolean algebras. Fundamenta Informaticae, 16(3–4):349–353, March 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gasarch:1992:LPE

[68] William I. Gasarch, Ramesh K. Sitaraman, Carl H. Smith, and Mahendran Velauthapillai. Learning programs with an easy to calculate set of errors. Fundamenta Informaticae, 16(3–4):355–370, March 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rauszer:1992:LIS

[69] Cecylia M. Rauszer. Logic for information systems. Fundamenta Informaticae, 16(3–4):371–382, March 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vakarelov:1992:IML

[70] Dimiter Vakarelov. Inductive modal logics. Fundamenta Informaticae, 16 (3–4):383–405, March 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Boutilier:1992:EEA

[71] Craig Boutilier. Epistemic entrenchment in autoepistemic logic. Fundamenta Informaticae, 17(1–2):5–29, September/October 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Eiter:1992:RPM

[72] Thomas Eiter and Georg Gottlob. Reasoning with parsimonious and moderately grounded expansions. Fundamenta Informaticae, 17(1–2):31–53, September/October 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fitting:1992:MVM

[73] Melvin Fitting. Many-valued modal logics II. Fundamenta Informaticae, 17(1–2):55–73, September/October 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lakemeyer:1992:PIQ

[74] Gerhard Lakemeyer. On perfect introspection with quantifying-in. Fundamenta Informaticae, 17(1–2):75–98, September/October 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Marek:1992:MMA

[75] V. Wiktor Marek and Mirosław Truszczyński. More on modal aspects of default logic. Fundamenta Informaticae, 17(1–2):99–116, September/ October 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Niemela:1992:DCA

[76] Ilkka N. F. Niemelä. On the decidability and complexity of autoepistemic reasoning. Fundamenta Informaticae, 17(1–2):117–155, September/ October 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Schwarz:1992:RAL

[77] Grigori Schwarz. Reflexive autoepistemic logic. Fundamenta Informaticae, 17(1–2):157–173, September/October 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

vanBenthem:1992:LP

[78] Johan van Benthem. Logic as programming. Fundamenta Informaticae, 17 (4):285–317, December 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Aceto:1992:HPC

[79] Luca Aceto. History preserving, causal and mixed-ordering equivalence over stable event structures (note). Fundamenta Informaticae, 17(4):319– 331, December 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Akdag:1992:QTU

[80] Herman Akdag, Michel De Glas, and Daniel Pacholczyk. A qualitative theory of uncertainty. Fundamenta Informaticae, 17(4):333–362, December 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Makinen:1992:GID

[81] Erkki Mäkinen. On grammatical inference and derivational complexity of context-free grammars. Fundamenta Informaticae, 17(4):363–368, December 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Aceto:1992:RDT

[82] Luca Aceto. Relating distributed, temporal and causal observations of simple processes. Fundamenta Informaticae, 17(4):369–397, December 1992. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Breazu-Tannen:1993:CET

[83] Val Breazu-Tannen and Albert R. Meyer. Conservativity of equational theories in typed lambda calculi. Fundamenta Informaticae, 19(1–2):1–49, September/October 1993. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Curien:1993:SI

[84] P.-L. Curien. Substitution up to isomorphism. Fundamenta Informaticae, 19(1–2):51–85, September/October 1993. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Giannini:1993:TIS

[85] Paola Giannini, Furio Honsell, and Simona Ronchi Della Rocca. Type inference: some results, some problems. Fundamenta Informaticae, 19(1– 2):87–125, September/October 1993. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jategaonkar:1993:TIE

[86] Lalita A. Jategaonkar and John C. Mitchell. Type inference with extended pattern matching and subtypes. Fundamenta Informaticae, 19 (1–2):127–165, September/October 1993. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Leivant:1993:LCC

[87] Daniel Leivant and Jean-Yves Marion. Lambda calculus characterizations of poly-time. Fundamenta Informaticae, 19(1–2):167–184, September/October 1993. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pfenning:1993:UPP

[88] Frank Pfenning. On the undecidability of partial polymorphic type reconstruction. Fundamenta Informaticae, 19(1–2):185–199, September/ October 1993. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Urzyczyn:1993:PRE

[89] Pawel Urzyczyn. Primitive recursion with existential types. Fundamenta Informaticae, 19(1–2):201–222, September/October 1993. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sakai:1993:FLP

[90] Hiroshi Sakai. On a framework for logic programming with incomplete information. Fundamenta Informaticae, 19(3–4):223–234, November/December 1993. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rauszer:1993:DRD

[91] Cecylia M. Rauszer. Dependencies in relational databases algebraic and logical approach. Fundamenta Informaticae, 19(3–4):235–274, November/ December 1993. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Biela:1993:RSD

[92] Andrzej Biela. Retrieval system and dynamic algorithm looking for axioms of notions defined by programs. Fundamenta Informaticae, 19(3–4):275– 301, November/December 1993. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Levene:1993:FPN

[93] Mark Levene and George Loizou. A fully precise null extended nested relational algebra. Fundamenta Informaticae, 19(3–4):303–342, November/ December 1993. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Novotny:1993:DWC

[94] Jiři Novotný and Miroslav Novotný. On dependence in Wille’s contexts. Fundamenta Informaticae, 19(3–4):343–353, November/December 1993. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kari:1993:DS

[95] Lila Kari, Alexandru Mateescu, Arto Salomaa, and Gheorghe Păun. Deletion sets. Fundamenta Informaticae, 19(3–4):355–370, November/ December 1993. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Marongiu:1993:UFT

[96] Gabriele Marongiu and Sauro Tulipani. Undecidable fragments of term algebras with subterm relation. Fundamenta Informaticae, 19(3–4):371– 382, November/December 1993. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Amir:1993:SP

[97] Amihood Amir and Carl H. Smith. The syntax of parallelism. Fundamenta Informaticae, 19(3–4):383–402, November/December 1993. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Murphy:1993:TDN

[98] David Murphy. Time and duration in noninterleaving concurrency. Fundamenta Informaticae, 19(3–4):403–416, November/December 1993. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Danko:1993:SPA

[99] Wiktor Dańko. The set of probabilistic algorithmic formulas valid in a finite structure is decidable with respect to its diagram. Fundamenta Informaticae, 19(3–4):417–431, November/December 1993. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rasiowa:1994:P

[100] Helena Rasiowa. Preface. Fundamenta Informaticae, 20(1–3):1–2, March/ April/May 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

deBakker:1994:BSC

[101] J. W. de Bakker and E. P. de Vink. Bisimulation semantics for concurrency with atomicity and action refinement. Fundamenta Informaticae, 20(1– 3):3–34, March/April/May 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Doyle:1994:RAR

[102] Jon Doyle. Reasoned assumptions and rational psychology. Fundamenta Informaticae, 20(1–3):35–73, March/April/May 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ehrenfeucht:1994:SS

[103] A. Ehrenfeucht and G. Rozenberg. Square systems. Fundamenta Informaticae, 20(1–3):75–111, March/April/May 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fitting:1994:KTV

[104] Melvin Fitting. Kleene’s three valued logics and their children. Fundamenta Informaticae, 20(1–3):113–131, March/April/May 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Harju:1994:RPS

[105] T. Harju and G. Rozenberg. Reductions for primitive 2-structures. Fundamenta Informaticae, 20(1–3):133–144, March/April/May 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Minker:1994:SDL

[106] Jack Minker and Carolina Ruiz. Semantics for disjunctive logic programs with explicit and default negation. Fundamenta Informaticae, 20(1–3): 145–192, March/April/May 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Montanari:1994:CCP

[107] Ugo Montanari and Daniel Yankelevich. Combining CCS and Petri nets via structural axioms. Fundamenta Informaticae, 20(1–3):193–229, March/April/May 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Trakhtenbrot:1994:CPN

[108] B. A. Trakhtenbrot. Compositional proofs for networks of processes. Fundamenta Informaticae, 20(1–3):231–275, March/April/May 1994. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tendera:1994:NAP

[109] Lidia Tendera. A note on asymptotic probabilities of existential secondorder minimal classes — the last step. Fundamenta Informaticae, 20(4): 277–285, December 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Noll:1994:TPS

[110] Thomas Noll and Heiko Vogler. Top-down parsing with simultaneous evaluation of noncircular attribute grammars. Fundamenta Informaticae, 20 (4):285–332, December 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Laneve:1994:DEC

[111] Cosimo Laneve. Distributive evaluations of λ-calculus. Fundamenta Informaticae, 20(4):333–352, December 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Baaz:1994:SPC

[112] Matthias Baaz and Alexander Leitsch. On Skolemization and proof complexity. Fundamenta Informaticae, 20(4):353–379, December 1994. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skandalis:1994:CIS

[113] Kostas Skandalis. On the characterization of input sets as output sets. Fundamenta Informaticae, 20(4):381–392, December 1994. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Etherington:1994:ISI

[114] David Etherington, Henry Kautz, and Kurt Konolige. Introduction to the special issue. Fundamenta Informaticae, 21(1–2):1–5, April 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stalnaker:1994:WNC

[115] Robert Stalnaker. What is a nonmonotonic consequence relation? Fundamenta Informaticae, 21(1–2):7–21, April 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Boddy:1994:CEM

[116] Mark Boddy, Robert P. Goldman, Keiji Kanazawa, and Lynn Andrea Stein. A critical examination of model preference defaults. Fundamenta Informaticae, 21(1–2):23–38, April 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Besnard:1994:PWS

[117] Philippe Besnard and Torsten Schaub. Possible worlds semantics for default logics. Fundamenta Informaticae, 21(1–2):39–66, April 1994. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Przymusinska:1994:SDE

[118] Halina Przymusinska and Teodor Przymusinski. Stationary default extensions. Fundamenta Informaticae, 21(1–2):67–87, April 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Schwind:1994:MLH

[119] Camilla Schwind and Pierre Siegel. A modal logic for hypothesis theory. Fundamenta Informaticae, 21(1–2):89–101, April 1994. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cadoli:1994:AID

[120] Marco Cadoli and Marco Schaerf. Approximate inference in default logic and circumscription. Fundamenta Informaticae, 21(1–2):103–112, April 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bertossi:1994:CGM

[121] Leopoldo E. Bertossi and Raymond Reiter. Circumscription and generic mathematical objects. Fundamenta Informaticae, 21(1–2):113–123, April 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

McCarty:1994:DDR

[122] L. Thorne McCarty. Defeasible Deontic reasoning. Fundamenta Informaticae, 21(1–2):125–148, April 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ginsberg:1994:WDC

[123] Matthew L. Ginsberg and Hugh W. Holbrook. What defaults can do that hierarchies can’t. Fundamenta Informaticae, 21(1–2):149–159, April 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Komorowski:1994:GEP

[124] J. Komorowski and Z. W. Ras. Guest Editor’s preface — selected papers from the ISMIS’93 Symposium. Fundamenta Informaticae, 21(3):161–162, August 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Robertson:1994:EPR

[125] Dave Robertson, Jaume Agustí, Jane Hesketh, and Jordi Levy. Expressing program requirements using refinement lattices. Fundamenta Informaticae, 21(3):163–182, August 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Leone:1994:MIL

[126] Nicola Leone, Luigi Palopoli, and Massimo Romeo. Modifying intensional logic knowledge. Fundamenta Informaticae, 21(3):183–203, August 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chen:1994:LOK

[127] Jianhua Chen. The logic of only knowing as a unified framework for nonmonotonic reasoning. Fundamenta Informaticae, 21(3):205–220, August 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chu:1994:MFS

[128] Heng Chu and David A. Plaisted. Model finding in semantically guided instance-based theorem proving. Fundamenta Informaticae, 21(3):221– 235, August 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Murray:1994:ACI

[129] Neil V. Murray and Erik Rosenthal. Adapting classical inference techniques to multiple-valued logics using signed formulas. Fundamenta Informaticae, 21(3):237–253, August 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:1994:O

[130] Anonymous. Obituary. Fundamenta Informaticae, 21(4):255–256, December 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moser:1994:CSA

[131] L. E. Moser, P. M. Melliar-Smith, G. Kutty, and Y. S. Ramakrishna. Completeness and soundness of axiomatizations for temporal logics without next. Fundamenta Informaticae, 21(4):257–305, December 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Baral:1994:VSF

[132] Chitta Baral. Varying selection functions to relate conditional logics and preferential models. Fundamenta Informaticae, 21(4):307–319, December 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Duntsch:1994:RRA

[133] Ivo Düntsch. Rough relation algebras. Fundamenta Informaticae, 21(4): 321–331, December 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cholak:1994:CLS

[134] Peter Cholak and Howard A. Blair. The complexity of local stratification. Fundamenta Informaticae, 21(4):333–344, December 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hungar:1994:ESS

[135] Hardi Hungar. Expressibility of the semantics of sequential programs in first-order logic. Fundamenta Informaticae, 21(4):345–365, December 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bozapalidis:1994:RTS

[136] Symeon Bozapalidis. Representable tree series. Fundamenta Informaticae, 21(4):367–389, December 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moshkov:1994:OPD

[137] Mikhail Moshkov. Optimization problems for decision trees. Fundamenta Informaticae, 21(4):391–401, December 1994. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Revesz:1995:GEN

[138] György E. Révész. Guest Editor’s note. Fundamenta Informaticae, 22(1– 2):1–2, April 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Asperti:1995:LLC

[139] Andrea Asperti. Linear logic, comonads and optimal reductions. Fundamenta Informaticae, 22(1–2):3–22, April 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hofmann:1995:ACH

[140] K. H. Hofmann and M. W. Mislove. All compact Hausdorff lambda models are degenerate. Fundamenta Informaticae, 22(1–2):23–52, April 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lambek:1995:CEC

[141] J. Lambek. Cut elimination for classical bilinear logic. Fundamenta Informaticae, 22(1–2):53–67, April 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Longo:1995:PTD

[142] Giuseppe Longo. Parametric and type-dependent polymorphism. Fundamenta Informaticae, 22(1–2):69–92, April 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mislove:1995:ABC

[143] Michael W. Mislove and Frank J. Oles. Adjunctions between categories of domains. Fundamenta Informaticae, 22(1–2):93–116, April 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moggi:1995:SEL

[144] Eugenio Moggi. A semantics for evaluation logic. Fundamenta Informaticae, 22(1–2):117–152, April 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Revesz:1995:CCE

[145] György E. Révész. Categorical combinators with explicit products. Fundamenta Informaticae, 22(1–2):153–166, April 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stark:1995:ADN

[146] Eugene W. Stark. An algebra of dataflow networks. Fundamenta Informaticae, 22(1–2):167–185, April 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wagner:1995:FPS

[147] Eric G. Wagner, Wafaa Khalil, and R. F. C. Walters. Fix-point semantics for programs in distributive categories. Fundamenta Informaticae, 22(1– 2):187–202, April 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moshkov:1995:ADD

[148] Mikhail Moshkov. About the depth of decision trees computing Boolean functions. Fundamenta Informaticae, 22(3):203–215, August 1995. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dassow:1995:GCC

[149] Jürgen Dassow, Gheorghe Păun, and Sorina Vicolov. On the generative capacity of certain classes of cooperating grammar systems. Fundamenta Informaticae, 22(3):217–226, August 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dix:1995:CTSa

[150] Jürgen Dix. A classification theory of semantics of normal logic programs: I. Strong properties. Fundamenta Informaticae, 22(3):227–255, August 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dix:1995:CTSb

[151] Jürgen Dix. A classification theory of semantics of normal logic programs: II. Weak properties. Fundamenta Informaticae, 22(3):257–288, August 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kolpakov:1995:CGR

[152] Roman M. Kolpakov. On the complexity of generation of rational numbers by Boolean functions. Fundamenta Informaticae, 22(3):289–298, August 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Degano:1995:FP

[153] Pierpaolo Degano and Leonarda Raffoni. Fairness and priority. Fundamenta Informaticae, 22(3):299–306, August 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Edelsbrunner:1995:P

[154] Herbert Edelsbrunner and Jerzy W. Jaromczyk. Preface. Fundamenta Informaticae, 22(4):307, December 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Eppstein:1995:DLT

[155] David Eppstein, Gary L. Miller, and Shang-Hua Teng. A deterministic linear time algorithm for geometric separators and its applications. Fundamenta Informaticae, 22(4):309–329, December 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Icking:1995:CDF

[156] Christian Icking, Rolf Klein, Ngoc-Minh Lé, and Lihong Ma. Convex distance functions in 3-space are different. Fundamenta Informaticae, 22 (4):331–352, December 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kirkpatrick:1995:TPS

[157] David Kirkpatrick and Jack Snoeyink. Tentative prune-and-search for computing fixed-points with applications to geometric computation. Fundamenta Informaticae, 22(4):353–370, December 1995. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bern:1995:CT

[158] Marshall Bern. Compatible tetrahedralizations. Fundamenta Informaticae, 22(4):371–384, December 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alon:1995:LNC

[159] Noga Alon, Sridhar Rajagopalan, and Subhash Suri. Long non-crossing configurations in the plane. Fundamenta Informaticae, 22(4):385–394, December 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ehrenfeucht:1995:GCT

[160] Andrzej Ehrenfeucht, Joost Engelfriet, Paulien ten Pas, and Grzegorz Rozenberg. Grammatical codes of trees and terminally coded grammars.

Fundamenta Informaticae, 23(1):1–32, January 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Meyer:1995:DLB

[161] J.-J. Ch. Meyer and W. van der Hoek. A default logic based on epistemic states. Fundamenta Informaticae, 23(1):33–65, January 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gasarch:1995:LQT

[162] William I. Gasarch, Efim B. Kinber, Mark G. Pleszkoch, Carl H. Smith, and Thomas Zeugmann. Learning via queries with teams and anomalies. Fundamenta Informaticae, 23(1):67–89, January 1995. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stewart:1995:RSC

[163] Iain A. Stewart. Reachability in some classes of acyclic Petri nets. Fundamenta Informaticae, 23(1):91–100, January 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

deKogel:1995:RAE

[164] Eric de Kogel. Relational algebra and equational proofs. Fundamenta Informaticae, 23(1):101–121, January 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cadoli:1995:AID

[165] Marco Cadoli and Marco Schaerf. Approximate inference in default logic and circumscription. Fundamenta Informaticae, 23(1):123–143, January 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Perlis:1995:PCT

[166] Don Perlis. Papers on context: Theory and practice. Fundamenta Informaticae, 23(2–4):145–148, August 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Attardi:1995:FV

[167] Guiseppe Attardi and Maria Simi. A formalization of viewpoints. Fundamenta Informaticae, 23(2–4):149–173, August 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Besnard:1995:ACB

[168] Philippe Besnard and Torsten Schaub. An approach to context-based default reasoning. Fundamenta Informaticae, 23(2–4):175–223, August 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bhatnagar:1995:CHU

[169] Raj Bhatnagar. Context hypothesization using probabilistic knowledge. Fundamenta Informaticae, 23(2–4):225–246, August 1995. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bouquet:1995:RAT

[170] Paolo Bouquet and Fausto Giunchiglia. Reasoning about theory adequacy. a new solution to the qualification problem. Fundamenta Informaticae, 23 (2–4):247–262, August 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Buvac:1995:MC

[171] Saša Buvač, Vanja Buvač, and Ian A. Mason. Metamathematics of contexts. Fundamenta Informaticae, 23(2–4):263–301, August 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Harper:1995:MMK

[172] Mary P. Harper and Randall A. Helzerman. Managing multiple knowledge sources in constraint-based parsing of spoken language. Fundamenta Informaticae, 23(2–4):303–353, August 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Miller:1995:CSC

[173] Michael Miller. Context shifts and clashes in dialogues: an active logic perspective. Fundamenta Informaticae, 23(2–4):355–370, August 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nirkhe:1995:FRT

[174] Madhura Nirkhe and Sarit Kraus. Formal real-time imagination. Fundamenta Informaticae, 23(2–4):371–394, August 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Plaisted:1995:EP

[175] David A. Plaisted. Editor’s preface. Fundamenta Informaticae, 24(1–2): 1–2, April 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gramlich:1995:ARB

[176] Bernhard Gramlich. Abstract relations between restricted termination and confluence properties of rewrite systems. Fundamenta Informaticae, 24(1–2):3–23, April 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Martin:1995:DOS

[177] Ursula Martin. On the diversity of orderings on strings. Fundamenta Informaticae, 24(1–2):25–46, April 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Steinbach:1995:SOH

[178] Joachim Steinbach. Simplification orderings: History of results. Fundamenta Informaticae, 24(1–2):47–87, April 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zantema:1995:TTR

[179] H. Zantema. Termination of term rewriting by semantic labelling. Fundamenta Informaticae, 24(1–2):89–105, April 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhang:1995:CRA

[180] Hantao Zhang. Contextual rewriting in automated reasoning. Fundamenta Informaticae, 24(1–2):107–123, April 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bockmayr:1995:NSA

[181] Alexander Bockmayr, Stefan Krischer, and Andreas Werner. Narrowing strategies for arbitrary canonical rewrite systems. Fundamenta Informaticae, 24(1–2):125–155, April 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gilleron:1995:RTL

[182] Rémy Gilleron and Sophie Tison. Regular tree languages and rewrite systems. Fundamenta Informaticae, 24(1–2):157–175, April 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bonacina:1995:CDM

[183] Maria Paola Bonacina and Jieh Hsiang. The clause-diffusion methodology for distributed deduction. Fundamenta Informaticae, 24(1–2):177– 207, April 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jacobs:1995:PPS

[184] Bart Jacobs. Parameters and parametrization in specification, using distributive categories. Fundamenta Informaticae, 24(3):209–250, August 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mitrana:1995:SGS

[185] Victor Mitrana. Similarity in grammar systems. Fundamenta Informaticae, 24(3):251–257, August 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Heckmann:1995:LBD

[186] Reinhold Heckmann. Lower bag domains. Fundamenta Informaticae, 24 (3):259–281, August 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gabarro:1995:RPL

[187] J. Gabarró and M. J. Serna. Rational processes and linear systems in GSP. Fundamenta Informaticae, 24(3):283–302, August 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Felea:1995:FCI

[188] Victor Felea. On the family of conditional implicational dependencies. Fundamenta Informaticae, 24(3):303–312, August 1995. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kutty:1995:AIL

[189] G. Kutty, L. E. Moser, P. M. Melliar-Smith, Y. S. Ramakrishna, and L. K. Dillon. Axiomatizations of interval logics. Fundamenta Informaticae, 24 (4):313–331, December 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Magott:1995:PES

[190] Jan Magott. Performance evaluation of systems defined in specification language Estelle. Fundamenta Informaticae, 24(4):333–357, December 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Seda:1995:TSL

[191] Anthony Karel Seda. Topology and the semantics of logic programs. Fundamenta Informaticae, 24(4):359–386, December 1995. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ramakrishna:1995:SPL

[192] Y. S. Ramakrishna. On the satisfiability problem for Lamport’s propositional temporal logic of actions and some of its extensions. Fundamenta Informaticae, 24(4):387–405, December 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gasquet:1995:NIR

[193] Olivier Gasquet. A new incompleteness result in Kripke semantics. Fundamenta Informaticae, 24(4):407–415, December 1995. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bethke:1996:ECI

[194] Inge Bethke and Piet Rodenburg. Equational constructor induction. Fundamenta Informaticae, 25(1):1–15, January 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fernau:1996:GLF

[195] Henning Fernau. On grammar and language families. Fundamenta Informaticae, 25(1):17–34, January 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bondarenko:1996:APC

[196] V. A. Bondarenko and S. V. Yurov. About a polyhedron of cubic graphs. Fundamenta Informaticae, 25(1):35–38, January 1996. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bujosa:1996:SED

[197] Andrés Bujosa and Regino Criado. Syntactic elements of declarative programming: Symbolic linear equations. Fundamenta Informaticae, 25(1): 39–48, January 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Narendran:1996:UMA

[198] Paliath Narendran. Unification modulo ACI + 1 + 0. Fundamenta Informaticae, 25(1):49–57, January 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kaufmann:1996:QNU

[199] Susanne Kaufmann and Martin Kummer. On a quantitative notion of uniformity. Fundamenta Informaticae, 25(1):59–78, January 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Korczynsky:1996:NC

[200] Waldemar Korczyńsky. On a notion of concurrence. Fundamenta Informaticae, 25(1):79–98, January 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

DeLaHiguera:1996:STS

[201] Colin De La Higuera and Marie-Catherine Daniel-Vatonne. On sets of terms: a study of a generalisation relation and of its algorithmic properties. Fundamenta Informaticae, 25(2):99–121, February 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stephens:1996:CST

[202] R. Stephens and B. C. Thompson. Cartesian stream transformer composition. Fundamenta Informaticae, 25(2):123–174, February 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Makinen:1996:NGI

[203] Erkki Mäkinen. A note on the grammatical inference problem for even linear languages. Fundamenta Informaticae, 25(2):175–181, February 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Penczek:1996:ATL

[204] Wojciech Penczek. Axiomatizations of temporal logics on trace systems. Fundamenta Informaticae, 25(2):183–200, February 1996. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moshkov:1996:CAD

[205] Mikhail Moshkov. Comparitive analysis of deterministic and nondeterministic decision tree complexity. Global approach. Fundamenta Informaticae, 25(2):201–214, February 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rozenberg:1996:P

[206] Grzegorz Rozenberg and Arto Salomaa. Preface. Fundamenta Informaticae, 25(3–4):215, March 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Berstel:1996:TAT

[207] J. Berstel and L. Boasson. Towards an algebraic theory of context-free languages. Fundamenta Informaticae, 25(3–4):217–239, March 1996. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Diekert:1996:NMC

[208] Volkert Diekert and Anca Muscholl. A note on Métivier’s construction of asynchronous automata for triangulated graphs. Fundamenta Informaticae, 25(3–4):241–246, March 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gecseg:1996:DA

[209] F. Gécseg and H. Jürgensen. Dependence in algebras. Fundamenta Informaticae, 25(3–4):247–256, March 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Honkala:1996:PSG

[210] Juha Honkala and Werner Kuich. On a power series generalization of ETOL languages. Fundamenta Informaticae, 25(3–4):257–270, March 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hromkovic:1996:TLB

[211] Juraj Hromkovič, Jarkko Kari, Lila Kari, and Dana Pardubská. Two lower bounds on distributive generation of languages. Fundamenta Informaticae, 25(3–4):271–284, March 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Derencourt:1996:CPW

[212] D. Derencourt, J. Karhumäki, M. Latteux, and A. Terlutte. On computational power of weighted finite automata. Fundamenta Informaticae, 25 (3–4):285–293, March 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Drewes:1996:DGP

[213] F. Drewes and H.-J. Kreowski. (un-)decidability of geometric properties of pictures generated by collage grammars. Fundamenta Informaticae, 25 (3–4):295–325, March 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lange:1996:CSF

[214] Klaus-Jörn Lange. Complexity and structure in formal language theory. Fundamenta Informaticae, 25(3–4):327–352, March 1996. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hoogeboom:1996:TLA

[215] Hendrik Jan Hoogeboom and Paulienten Pas. Text languages in an algebraic framework. Fundamenta Informaticae, 25(3–4):353–380, March 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paun:1996:CGP

[216] Gheorghe Păun, Grzegorz Rozenberg, and Arto Salomaa. Contextual grammars: Parallelism and blocking of derivation. Fundamenta Informaticae, 25(3–4):381–397, March 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Giammarresi:1996:TDF

[217] Dora Giammarresi and Antonio Restivo. Two-dimensional finite state recognizability. Fundamenta Informaticae, 25(3–4):399–422, March 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Guo:1996:SL

[218] L. Guo, K. Salomaa, and S. Yu. On synchronization languages. Fundamenta Informaticae, 25(3–4):423–436, March 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Biswas:1996:VFS

[219] Ranjit Biswas. On I–V fuzzy subgroups. Fundamenta Informaticae, 26 (1):1–9, January 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Menezes:1996:SPN

[220] P. Blauth Menezes and J. Félix Costa. Synchronization in Petri nets. Fundamenta Informaticae, 26(1):11–22, January 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Spilker:1996:SUR

[221] Heike Spilker and Dietmar Wätjen. Some undecidability results concerning k-limited 0L systems. Fundamenta Informaticae, 26(1):23–30, January 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rishe:1996:UMN

[222] Naphtali Rishe. A universal model for non-procedural database languages. Fundamenta Informaticae, 26(1):31–57, January 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Boldi:1996:MCA

[223] Paolo Boldi. Maximal chains and antichains in strongly Noetherian semiorders. Fundamenta Informaticae, 26(1):59–80, January 1996. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Szalas:1996:NDF

[224] Andrzej Szalas. On natural deduction in first-order fixpoint logics. Fundamenta Informaticae, 26(1):81–94, January 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hirshfeld:1996:ULE

[225] Yoram Hirshfeld. Undecidability of language equivalence for generalized regular expressions. Fundamenta Informaticae, 26(1):95–102, January 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fokkink:1996:CAP

[226] Wan Fokkink. A complete axiomatization for prefix iteration in branching bisimulation. Fundamenta Informaticae, 26(2):103–113, May 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fu:1996:RMG

[227] Yuxi Fu. Recursive models of general inductive types. Fundamenta Informaticae, 26(2):115–131, May 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Knijnenburg:1996:NSP

[228] Peter M. W. Knijnenburg. A note on the Smyth powerdomain construction. Fundamenta Informaticae, 26(2):133–139, May 1996. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

vanBakel:1996:RIT

[229] Steffen van Bakel. Rank 2 intersection type assignment in term rewriting systems. Fundamenta Informaticae, 26(2):141–166, May 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Konikowska:1996:MAM

[230] Beata Konikowska. McCarthy algebras: a model of McCarthy’s logical calculus. Fundamenta Informaticae, 26(2):167–203, May 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Engels:1996:P

[231] G. Engels, H. Ehrig, and G. Rozenberg. Preface. Fundamenta Informaticae, 26(3–4):205, June 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ariola:1996:ETG

[232] Zena M. Ariola and Jan Willem Klop. Equational term graph rewriting. Fundamenta Informaticae, 26(3–4):207–240, June 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Corradini:1996:GP

[233] A. Corradini, U. Montanari, and F. Rossi. Graph processes. Fundamenta Informaticae, 26(3–4):241–265, June 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Drewes:1996:LBG

[234] Frank Drewes. A lower bound on the growth of functions computed by tree transducers. Fundamenta Informaticae, 26(3–4):267–285, June 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Habel:1996:GGN

[235] Annegret Habel, Reiko Heckel, and Gabriele Taentzer. Graph grammars with negative application conditions. Fundamenta Informaticae, 26(3–4): 287–313, June 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Janssens:1996:ASE

[236] D. Janssens and T. Mens. Abstract semantics for ESM systems. Fundamenta Informaticae, 26(3–4):315–339, June 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mosbah:1996:PGG

[237] Mohamed Mosbah. Probabilistic graph grammars. Fundamenta Informaticae, 26(3–4):341–362, June 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Schurr:1996:LBP

[238] Andy Schürr. Logic based programmed structure rewriting systems. Fundamenta Informaticae, 26(3–4):363–385, June 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Taentzer:1996:TSA

[239] Gabriele Taentzer. Towards synchronous and asynchronous graph transformations. Fundamenta Informaticae, 26(3–4):387–406, June 1996. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wagner:1996:DOB

[240] Annika Wagner and Martin Gogolla. Defining operational behavior of object specifications by attributed graph transformations. Fundamenta Informaticae, 26(3–4):407–431, June 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sopena:1996:EGR

[241] Eric Sopena. Expanding graph relabeling systems have the power of recursive enumerability. Fundamenta Informaticae, 27(1):1–25, July 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

LaTorre:1996:PWS

[242] Salvatore La Torre, Margherita Napoli, and Domenico Parente. Parallel word substitution. Fundamenta Informaticae, 27(1):27–36, July 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bergstra:1996:PME

[243] J. A. Bergstra and Gh. Ştefănescu. Processes with multiple entries and exits modulo isomorphism and modulo bisimulation. Fundamenta Informaticae, 27(1):37–56, July 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Liau:1996:AFR

[244] Churn-Jung Liau. An algebraic formalization of the relationship between evidential structures and data tables. Fundamenta Informaticae, 27(1): 57–76, July 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Marcus:1996:SSD

[245] Leo Marcus. Syntactic and semantic dependence of array-arithmetic sentences, with an application to program verification. Fundamenta Informaticae, 27(1):77–100, July 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skowron:1996:P

[246] Andrzej Skowron and Wojciech Ziarko. Preface. Fundamenta Informaticae, 27(2–3):101–102, August 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pawlak:1996:RSR

[247] Zdzislaw Pawlak. Rough sets, rough relations and rough functions. Fundamenta Informaticae, 27(2–3):103–108, August 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bryniarski:1996:FCR

[248] Edward Bryniarski. Formal conception of rough sets. Fundamenta Informaticae, 27(2–3):109–136, August 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lin:1996:FOR

[249] T. Y. Lin and Qing Liu. First-order rough logic I: Approximate reasoning via rough sets. Fundamenta Informaticae, 27(2–3):137–153, August 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Katzberg:1996:VPE

[250] Jack David Katzberg and Wojciech Ziarko. Variable precision extension of rough sets. Fundamenta Informaticae, 27(2–3):155–168, August 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kent:1996:RCA

[251] Robert E. Kent. Rough concept analysis: a synthesis of rough sets and formal concept analysis. Fundamenta Informaticae, 27(2–3):169–181, August 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kryszkiewicz:1996:CRC

[252] Marzena Kryszkiewicz and Henryk Rybinski. Computation of reducts of composed information systems. Fundamenta Informaticae, 27(2–3):183– 195, August 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moshkov:1996:SBM

[253] Mikhail Moshkov. Some bounds on minimal decision tree depth. Fundamenta Informaticae, 27(2–3):197–203, August 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pagliani:1996:RSN

[254] Piero Pagliani. Rough sets and Nelson algebras. Fundamenta Informaticae, 27(2–3):205–219, August 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pomykala:1996:NCI

[255] Janusz Pomykala and Erik De Haas. A note on categories of information systems. Fundamenta Informaticae, 27(2–3):221–227, August 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Slowinski:1996:RSR

[256] Roman Slowiński and Jerzy Stefanowski. Rough-set reasoning about uncertain data. Fundamenta Informaticae, 27(2–3):229–243, August 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skowron:1996:TAS

[257] Andrzej Skowron and Jaroslaw Stepaniuk. Tolerance approximation spaces. Fundamenta Informaticae, 27(2–3):245–253, August 1996. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skowron:1996:AMM

[258] Andrzej Skowron and Lech Polkowski. Analytical morphology: Mathematical morphology of decision tables. Fundamenta Informaticae, 27 (2–3):255–271, August 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tsumoto:1996:CAF

[259] Shusaku Tsumoto and Hiroshi Tanaka. A common algebraic framework of empirical learning methods based on rough sets and matroid theory. Fundamenta Informaticae, 27(2–3):273–288, August 1996. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yao:1996:CRS

[260] Y. Y. Yao and Xining Li. Comparison of rough-set and interval-set models for uncertain reasoning. Fundamenta Informaticae, 27(2–3):289–298, August 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zytkow:1996:ADE

[261] Jan M. Zytkow. Automated discovery of empirical laws. Fundamenta Informaticae, 27(2–3):299–318, August 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Levene:1996:CCD

[262] Mark Levene and George Loizou. Categorisation of computable database queries. Fundamenta Informaticae, 27(4):319–348, September 1996. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Corradini:1996:FPO

[263] Flavio Corradini and Rocco De Nicola. On four partial ordering semantics for a process calculus. Fundamenta Informaticae, 27(4):349–383, September 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lobo:1996:ACO

[264] Jorge Lobo and Carlos Uzcátegui. Abductive change operators. Fundamenta Informaticae, 27(4):385–411, September 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Makinen:1996:TUA

[265] Erkki Mäkinen. (0, 1)-totality is undecidable for arbitrary context-free grammars. Fundamenta Informaticae, 27(4):413–415, September 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Marek:1996:Pa

[266] Wiktor Marek, Lech Polkowski, and Andrzej Skowron. Preface. Fundamenta Informaticae, 28(1–2):??, November 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alferes:1996:BRN

[267] José Júlio Alferes, Luís Moniz Pereira, and Teodor C. Przymusinski. Belief revision in non-monotonic reasoning and logic programming. Fundamenta Informaticae, 28(1–2):1–22, November 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Apt:1996:MVL

[268] Krzysztof R. Apt and Rachel Ben-Eliyahu. Meta-variables in logic programming, or in praise of ambivalent syntax. Fundamenta Informaticae, 28(1–2):23–36, November 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Blair:1996:DD

[269] Howard A. Blair. Designing dependencies. Fundamenta Informaticae, 28 (1–2):37–54, November 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Blikle:1996:WD

[270] Andrzej Blikle. Why denotational? Fundamenta Informaticae, 28(1– 2):55–85, November 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dix:1996:RDN

[271] Jürgen Dix, Georg Gottlob, and Wiktor Marek. Reducing disjunctive to non-disjunctive semantics by shift-operations. Fundamenta Informaticae, 28(1–2):87–100, November 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fitting:1996:MHT

[272] Melvin Fitting. A modal Herbrand theorem. Fundamenta Informaticae, 28(1–2):101–122, November 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gottlob:1996:ASM

[273] Georg Gottlob and Mirosław Truszczyński. Approximating the stable model semantics is hard. Fundamenta Informaticae, 28(1–2):123–128, November 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hsiang:1996:AAP

[274] Jieh Hsiang and Anita Wasilewska. Automating algebraic proofs in algebraic logic. Fundamenta Informaticae, 28(1–2):129–140, November 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mirkowska:1996:ASD

[275] Grazyna Mirkowska and Andrzej Salwicki. The algebraic specifications do not have the Tennenbaum property. Fundamenta Informaticae, 28(1–2): 141–152, November 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Plaza:1996:LPP

[276] Jan A. Plaza. Logic programming from the perspective of algebraic semantics. Fundamenta Informaticae, 28(1–2):153–164, November 1996. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pratt:1996:SIP

[277] Vaughan Pratt and Jerzy Tiuryn. Satisfiability of inequalities in a poset. Fundamenta Informaticae, 28(1–2):165–182, November 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Trakhtenbrot:1996:PCP

[278] B. A. Trakhtenbrot. On the power of compositional proofs for nets: Relationships between completeness and modularity. Fundamenta Informaticae, 28(1–2):183–195, November 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Urzyczyn:1996:PRT

[279] Paweł Urzyczyn. Positive recursive type assignment. Fundamenta Informaticae, 28(1–2):197–209, November 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Marek:1996:Pb

[280] Wiktor Marek, Lech Polkowski, and Andrzej Skowron. Preface. Fundamenta Informaticae, 28(3–4):??, December 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Banerjee:1996:RST

[281] Mohua Banerjee and Mihir K. Chakraborty. Rough sets through algebraic logic. Fundamenta Informaticae, 28(3–4):211–221, December 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bochman:1996:LBN

[282] Alexander Bochman. On a logical basis of normal logic programs. Fundamenta Informaticae, 28(3–4):223–245, December 1996. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ho:1996:MLR

[283] Nguyen Cat Ho. A method in linguistic reasoning on a knowledge base representing by sentences with linguistic belief degree. Fundamenta Informaticae, 28(3–4):247–259, December 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Doherty:1996:RRC

[284] Patrick Doherty, Witold Łukaszewicz, and Andrzej Szałas. A reduction result for circumscribed semi-Horn formulas. Fundamenta Informaticae, 28(3–4):261–271, December 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Green:1996:RSM

[285] John Green, Neil Horne, Ewa Orłowska, and Paul Siemens. A rough set model of information retrieval. Fundamenta Informaticae, 28(3–4): 273–296, December 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grigoriev:1996:SPN

[286] Dima Grigoriev, Marek Karpinski, and Andrew M. Odlyzko. Short proofs for nondivisibility of sparse polynomials under the extended Riemann hypothesis. Fundamenta Informaticae, 28(3–4):297–301, December 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gomolinska:1996:SNM

[287] Anna Gomolińska. On semantics for the nonmonotonic modal formalization of the logic of acceptance and rejection. Fundamenta Informaticae, 28(3–4):303–313, December 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paun:1996:PCG

[288] Gheorghe Păun, Lech Polkowski, and Andrzej Skowron. Parallel communicating grammar systems with negotiation. Fundamenta Informaticae, 28(3–4):315–330, December 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Radev:1996:AS

[289] Slavian Radev. Argumentation systems. Fundamenta Informaticae, 28(3– 4):331–346, December 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skarbek:1996:DBS

[290] Władysław Skarbek. Dynamic behaviour of spatial signal OR-graphs. Fundamenta Informaticae, 28(3–4):347–352, December 1996. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Suraj:1996:DCD

[291] Zbigniew Suraj. Discovery of concurrent data models from experimental tables: A rough set approach. Fundamenta Informaticae, 28(3–4):353–376, December 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Schlechta:1996:TSA

[292] Karl Schlechta. A two-stage approach to first order default reasoning. Fundamenta Informaticae, 28(3–4):377–402, December 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Winkowski:1996:CWP

[293] Józef Winkowski. Concatenable weighted pomsets and their applications to modelling processes of Petri nets. Fundamenta Informaticae, 28(3–4): 403–421, December 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wroblewski:1996:TFO

[294] Jakub Wróblewski. Theoretical foundations of order-based genetic algorithms. Fundamenta Informaticae, 28(3–4):423–430, December 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hotz:1997:USD

[295] Günter Hotz and Gisela Pitsch. On using semi-Dyck sets to analyse coupled-context-free languages. Fundamenta Informaticae, 29(1–2):1–26, January 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dembinski:1997:STC

[296] Piotr Dembiński. Semantics of timed concurrent systems. Fundamenta Informaticae, 29(1–2):27–50, January 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Baeten:1997:DTP

[297] J. C. M. Baeten and J. A. Bergstra. Discrete time process algebra: Absolute time, relative time and parametric time. Fundamenta Informaticae, 29(1–2):51–76, January 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Huzar:1997:SSR

[298] Zbigniew Huzar and Jan Magott. Syntax and semantics of real-time and performance evaluation extensions of specification language lotos. Fundamenta Informaticae, 29(1–2):77–96, January 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Seda:1997:QMS

[299] Anthony Karel Seda. Quasi-metrics and the semantics of logic programs. Fundamenta Informaticae, 29(1–2):97–117, January 1997. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Duntsch:1997:AAA

[300] Ivo Düntsch and Günther Gediga. Algebraic aspects of attribute dependencies in information systems. Fundamenta Informaticae, 29(1–2):119– 133, January 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Peters:1997:MRT

[301] James F. Peters. Mechanization of real-time linear CSP with higher order logic. Fundamenta Informaticae, 29(1–2):135–163, January 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Egly:1997:DTN

[302] Uwe Egly. On definitional transformations to normal form for intuitionistic logic. Fundamenta Informaticae, 29(1–2):165–201, January 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alexandre:1997:MCL

[303] Francis Alexandre and Khaled Bsaïes. A methodology for constructing logic programs. Fundamenta Informaticae, 29(3):203–223, February 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Baier:1997:HIE

[304] Christel Baier and Mila Majster-Cederbaum. How to interpret and establish consistency results for semantics of concurrent programming languages. Fundamenta Informaticae, 29(3):225–256, February 1997. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mohnen:1997:CCP

[305] Markus Mohnen. A compiler correctness proof for the static link technique by means of evolving algebras. Fundamenta Informaticae, 29(3):257–303, February 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ehrenfeucht:1997:SNT

[306] A. Ehrenfeucht, G. Rozenberg, and K. Salomaa. Semantics of nonsequential tree-based computation schemes. Fundamenta Informaticae, 29 (4):305–324, March 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fachini:1997:CHS

[307] Emanuela Fachini and Angelo Monti. Chomsky hierarchy and systolic Y -tree automata. Fundamenta Informaticae, 29(4):325–339, March 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ferrari:1997:ACC

[308] Gian-Luigi Ferrari. Atomicity and concurrency control in process calculi. Fundamenta Informaticae, 29(4):341–368, March 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jarvinen:1997:RDS

[309] Jouni Järvinen. A representation of dependence spaces and some basic algorithms. Fundamenta Informaticae, 29(4):369–382, March 1997. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kundu:1997:NMC

[310] Sukhamay Kundu and Jianhua Chen. A new method of circumscribing beliefs the propositional case. Fundamenta Informaticae, 29(4):383–390, March 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gasarch:1997:AQV

[311] William Gasarch and Mahendran Velauthapillai. Asking questions versus verifiability. Fundamenta Informaticae, 30(1):1–9, January 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kiehn:1997:DNI

[312] Astrid Kiehn and Matthew Hennessy. On the decidability of noninterleaving process equivalences. Fundamenta Informaticae, 30(1):11–30, January 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Madlener:1997:SUR

[313] Klaus Madlener and Friedrich Otto. Some undecidability results for finitely generated Thue congruences on a two-letter alphabet. Fundamenta Informaticae, 30(1):31–44, January 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paun:1997:CSC

[314] Gheorghe Păun. Controlled H systems and Chomsky hierarchy. Fundamenta Informaticae, 30(1):45–57, January 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Peltier:1997:TAA

[315] Nicolas Peltier. Tree automata and automated model building. Fundamenta Informaticae, 30(1):59–81, January 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Trakhtenbrot:1997:PCP

[316] B. A. Trakhtenbrot. On the power of compositional proofs for nets: Relationships between completenesess and modularity. Fundamenta Informaticae, 30(1):83–95, January 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Winkowski:1997:RPP

[317] Józef Winkowski. A representation of processes of Petri nets by matrices. Fundamenta Informaticae, 30(1):97–107, January 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhang:1997:DIC

[318] Yan Zhang and Norman Y. Foo. Deriving invariants and constraints from action theories. Fundamenta Informaticae, 30(1):109–123, January 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Basarab:1997:DFG

[319] Şerban A. Basarab. Directions and foldings on generalized trees. Fundamenta Informaticae, 30(2):125–149, May 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Findler:1997:STR

[320] Nicholas V. Findler. Some techniques of reducing the dangers of combinatorial explosion in automatic knowledge acquisition. Fundamenta Informaticae, 30(2):151–160, May 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ilie:1997:CCM

[321] Lucian Ilie. On the computational complexity of Marcus contextual languages. Fundamenta Informaticae, 30(2):161–167, May 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kandulski:1997:GAL

[322] Maciej Kandulski. On generalized Ajdukiewicz and Lambek calculi and grammars. Fundamenta Informaticae, 30(2):169–181, May 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Makinen:1997:PAP

[323] Erkki Mäkinen and Ferucio Laurenţiu Ţiplea. Pattern ambiguities for pure context — free grammars. Fundamenta Informaticae, 30(2):183–191, May 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Llabres:1997:DPH

[324] Mercè Llabrés and Francesc Rosselló. Double-pushout hypergraph rewriting using total conformisms. Fundamenta Informaticae, 30(2):193–226, May 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Michalewicz:1997:GEN

[325] Maciej Michalewicz and Zbigniew W. Ras. Guest Editors’ note. Fundamenta Informaticae, 30(3–4):??, June 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Clark:1997:IIC

[326] Peter Clark, Cao Feng, Stan Matwin, and Ko Fung. Improving image classification by combining statistical, case-based and model-based prediction methods. Fundamenta Informaticae, 30(3–4):227–240, June 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Druzdzel:1997:FUP

[327] Marek J. Druzdzel. Five useful properties of probabilistic knowledge representations from the point of view of intelligent systems. Fundamenta Informaticae, 30(3–4):241–254, June 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grzymala-Busse:1997:CUE

[328] Jerzy W. Grzymala-Busse. Classification of unseen examples under uncertainty. Fundamenta Informaticae, 30(3–4):255–267, June 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kersten:1997:FAS

[329] Gregory E. Kersten and Stan Szpakowicz. Formal account of sequential decision-making in a co-operative setting. Fundamenta Informaticae, 30 (3–4):269–281, June 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Michalski:1997:SKD

[330] Ryszard S. Michalski. Seeking knowledge in the deluge of facts. Fundamenta Informaticae, 30(3–4):283–297, June 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mrozek:1997:KRF

[331] Adam Mrózek and Leszek Płonka. Knowledge representation in fuzzy and rough controllers. Fundamenta Informaticae, 30(3–4):299–311, June 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ras:1997:QAA

[332] Zbigniew W. Ras and Sucheta Joshi. Query approximate answering system for an incomplete DKBS. Fundamenta Informaticae, 30(3–4):313– 324, June 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rusinkiewicz:1997:TWD

[333] Marek Rusinkiewicz and Mauro Bregolin. Transactional workflows in distributed systems. Fundamenta Informaticae, 30(3–4):325–344, June 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skowron:1997:DAS

[334] Andrzej Skowron and Lech Polkowski. Decision algorithms: a survey of rough set — theoretic methods. Fundamenta Informaticae, 30(3–4):345– 358, June 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wierzchon:1997:RFE

[335] S. T. Wierzchoń, M. A. Kłopotek, and M. Michalewicz. Reasoning and facts explanation in valuation based systems. Fundamenta Informaticae, 30(3–4):359–371, June 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ziarko:1997:MLT

[336] Wojciech Ziarko and Ning Shan. Machine learning through data classification and reduction. Fundamenta Informaticae, 30(3–4):373–382, June 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zytkow:1997:CTF

[337] Jan M. Żytkow and Robert Zembowicz. Contingency tables as the foundation for concepts, concept hierarchies, and rules: The 49er system approach. Fundamenta Informaticae, 30(3–4):383–399, June 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Buszkowski:1997:RRS

[338] Wojciech Buszkowski and Mirosława Kołowska-Gawiejnowicz. Representation of residuated semigroups in some algebras of relations (the method of canonical models). Fundamenta Informaticae, 31(1):1–12, July 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Esparza:1997:PNC

[339] Javier Esparza. Petri nets, commutative context-free grammars, and basic parallel processes. Fundamenta Informaticae, 31(1):13–25, July 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grzymala-Busse:1997:NVR

[340] Jerzy W. Grzymala-Busse. A new version of the rule induction system LERS. Fundamenta Informaticae, 31(1):27–39, July 1997. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Koshiba:1997:HSL

[341] Takeshi Koshiba. On a hierarchy of slender languages based on control sets. Fundamenta Informaticae, 31(1):41–47, July 1997. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Michalski:1997:LDS

[342] Ryszard S. Michalski and Ibrahim F. Imam. On learning decision structures. Fundamenta Informaticae, 31(1):49–64, July 1997. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Romanowska:1997:DSA

[343] Anna Romanowska. From dissemilattices to snack algebras. Fundamenta Informaticae, 31(1):65–77, July 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lee:1997:FVD

[344] Shie-Jue Lee and Wei-Jer Lin. Formal verification and diagnosis of combinational circuit designs with propositional logic. Fundamenta Informaticae, 31(1):79–105, July 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Majorinc:1997:ERN

[345] Kazimir Majorinc. Extension rule for non-clausal propositional calculus. Fundamenta Informaticae, 31(2):107–116, August 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Martin-Vide:1997:NCN

[346] Carlos Martín-Vide. Natural computation for natural language. Fundamenta Informaticae, 31(2):117–124, August 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Matskin:1997:PSS

[347] Mihhail Matskin and Jan Komorowski. Partial structural synthesis of programs. Fundamenta Informaticae, 31(2):125–144, August 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moshkov:1997:BAW

[348] Mikhail Moshkov and Igor Chikalov. Bounds on average weighted depth of decision trees. Fundamenta Informaticae, 31(2):145–156, August 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moshkov:1997:UUB

[349] Mikhail Moshkov. Unimprovable upper bounds on time complexity of decision trees. Fundamenta Informaticae, 31(2):157–184, August 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paun:1997:CSS

[350] Andrei Păun. Controlled H systems of small radius. Fundamenta Informaticae, 31(2):185–193, August 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skordev:1997:AAS

[351] Dimiter Skordev. An abstract approach to some loop detection problems. Fundamenta Informaticae, 31(2):195–212, August 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Banerjee:1997:RSV

[352] Mohua Banerjee. Rough sets and 3-valued Łukasiewicz logic. Fundamenta Informaticae, 31(3–4):213–220, September 1997. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Brewka:1997:WFS

[353] Gerhard Brewka and Georg Gottlob. Well-founded semantics for default logic. Fundamenta Informaticae, 31(3–4):221–236, September 1997. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Burkhard:1997:DBA

[354] Hans-Dieter Burkhard. Defining BDI with abstract languages. Fundamenta Informaticae, 31(3–4):237–252, September 1997. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chitil:1997:SCS

[355] Olaf Chitil. The ζ-semantics: A comprehensive semantics for functional programs. Fundamenta Informaticae, 31(3–4):253–294, September 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Deneva:1997:MLL

[356] Ana Deneva and Dimiter Vakarelov. Modal logics for local and global similarity relations. Fundamenta Informaticae, 31(3–4):295–304, September 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maggiolo-Schettini:1997:PCE

[357] Andrea Maggiolo-Schettini and Gionata Matteucci. Processes in causeeffect systems. Fundamenta Informaticae, 31(3–4):305–335, September 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maggiolo-Schettini:1997:POR

[358] Andrea Maggiolo-Schettini and Józef Winkowski. A partial order representation of processes of transforming graphs. Fundamenta Informaticae, 31(3–4):337–355, September 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Montanari:1997:CFS

[359] Ugo Montanari and Gioia Ristori. A concurrent functional semantics for a process algebra based on action systems. Fundamenta Informaticae, 31(3–4):357–377, September 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rey:1997:BPC

[360] Jean-François Rey. The block product of categories and Tilson’s division. Fundamenta Informaticae, 31(3–4):379–400, September 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rey:1997:KRM

[361] Jean-François Rey. The kernel of a relational morphism of categories. Fundamenta Informaticae, 31(3–4):401–421, September 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Frias:1997:FAA

[362] Marcelo F. Frias, Gabriel A. Baum, and Armando M. Haeberer. Fork algebras in algebra, logic and computer science. Fundamenta Informaticae, 32(1):1–25, October 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kasangian:1997:TLT

[363] Stefano Kasangian and Sebastiano Vigna. The topos of labelled trees: A categorical semantics for SCCS. Fundamenta Informaticae, 32(1):27–45, October 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pigozzi:1997:LAA

[364] Don Pigozzi and Antonino Salibra. Lambda abstraction algebras: Coordinatizing models of Lambda calculus. Fundamenta Informaticae, 32 (1):47–90, October 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tyszkiewicz:1997:QAC

[365] Jerzy Tyszkiewicz. Queries and algorithms computable by polynomial time existential reflective machines. Fundamenta Informaticae, 32(1):91– 105, October 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Komorowski:1997:GEP

[366] J. Komorowski and Z. W. Raś. Guest Editors’ preface — in memory of Cecylia Rauszer. Fundamenta Informaticae, 32(2):??, November 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chakraborty:1997:SPS

[367] Mihir K. Chakraborty and Ewa Orlowska. Substitutivity principles in some theories of uncertainty. Fundamenta Informaticae, 32(2):107–120, November 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gaasterland:1997:QAA

[368] Terry Gaasterland and Jorge Lobo. Qualifying answers according to user needs and preferences. Fundamenta Informaticae, 32(2):121–137, November 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Marek:1997:CRN

[369] V. Wiktor Marek, Anil Nerode, and Jeffrey B. Remmel. Complexity of recursive normal default logic. Fundamenta Informaticae, 32(2):139–147, November 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paun:1997:RSA

[370] Gheorghe Păun, Lech Polkowski, and Andrzej Skowron. Rough set approximations of languages. Fundamenta Informaticae, 32(2):149–162, November 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ruiz:1997:CCW

[371] Carolina Ruiz and Jack Minker. Combining closed world assumptions with stable negation. Fundamenta Informaticae, 32(2):163–181, November 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Segerberg:1997:PTB

[372] Krister Segerberg. Proposal for a theory of belief revision along the lines of Lindström and Rabinowicz. Fundamenta Informaticae, 32(2):183–191, November 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alessi:1997:CPL

[373] Fabio Alessi, Mariangiola Dezani-Ciancaglini, and Ugo de’ Liguoro. A convex powerdomain over lattices: its logic and λ-calculus. Fundamenta Informaticae, 32(3–4):193–250, December 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bucciarelli:1997:BMR

[374] Antonio Bucciarelli. Bi-models: Relational versus domain-theoretic approaches. Fundamenta Informaticae, 32(3–4):251–266, December 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bujosa:1997:UNS

[375] Andrés Bujosa, Regino Criado, and Miguel A. Hernandez-Medina. Unification: Nothing but the solution of a system of linear equations. Fundamenta Informaticae, 32(3–4):267–280, December 1997. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Castilho:1997:MTP

[376] Marcos A. Castilho, Luis Fariñas del Cerro, Olivier Gasquet, and Andreas Herzig. Modal tableaux with propagation rules and structural rules. Fundamenta Informaticae, 32(3–4):281–297, December 1997. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chakraborty:1997:GCS

[377] Mihir K. Chakraborty and Sanjukta Basu. Graded consequence and some metalogical notions generalized. Fundamenta Informaticae, 32(3–4):299– 311, December 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chlebus:1997:TOT

[378] Bogdan S. Chlebus, Krzysztof Diks, and Andrzej Pelc. Transition-optimal token distribution. Fundamenta Informaticae, 32(3–4):313–328, December 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fokkink:1997:ARP

[379] Wan Fokkink. An axiomatization for regular processes in timed branching bisimulation. Fundamenta Informaticae, 32(3–4):329–340, December 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Honkala:1997:DPC

[380] Juha Honkala. Decision problems concerning a power series generalization of DTOL systems. Fundamenta Informaticae, 32(3–4):341–348, December 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Khamsi:1997:DSL

[381] Mohamed A. Khamsi and Driss Misane. Disjunctive signed logic programs. Fundamenta Informaticae, 32(3–4):349–357, December 1997. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skarbek:1997:RIC

[382] Władysław Skarbek. On relation of image compression and image association. Fundamenta Informaticae, 32(3–4):359–371, December 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tiplea:1997:JPN

[383] Ferucio Laurenţiu Ţiplea and Erkki Mäkinen. Jumping Petri nets. Specific properties. Fundamenta Informaticae, 32(3–4):373–392, December 1997. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:1998:MMT

[384] Ludwik Czaja. Minimal-maximal time cause-effect structures. Fundamenta Informaticae, 33(1):1–16, January 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:1998:CES

[385] Ludwik Czaja. Cause-effect structures — structural and semantic properties revisited. Fundamenta Informaticae, 33(1):17–42, January 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dimov:1998:SCS

[386] Georgi Dimov and Dimiter Vakarelov. On Scott consequence systems. Fundamenta Informaticae, 33(1):43–70, January 1998. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Schmitt:1998:CS

[387] Vincent Schmitt. CR-structures. Fundamenta Informaticae, 33(1):71–83, January 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vigneron:1998:ADT

[388] Laurent Vigneron. Automated deduction techniques for studying rough algebras. Fundamenta Informaticae, 33(1):85–103, January 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Calude:1998:CAR

[389] Cristian S. Calude and Peter H. Hertling. Computable approximations of reals: An information-theoretic analysis. Fundamenta Informaticae, 33 (2):105–120, February 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pankowski:1998:PPR

[390] Tadeusz Pankowski. Powerdomain of paths for representing object structures. Fundamenta Informaticae, 33(2):121–148, February 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pigozzi:1998:LAA

[391] Don Pigozzi and Antonino Salibra. Lambda abstraction algebras: Coordinatizing models of Lambda calculus. Fundamenta Informaticae, 33(2): 149–200, February 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Plump:1998:TGR

[392] Detlef Plump. Termination of graph rewriting is undecidable. Fundamenta Informaticae, 33(2):201–209, February 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Marcus:1998:P

[393] Solomon Marcus. Preface. Fundamenta Informaticae, 33(3):??, March 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gramatovici:1998:EPC

[394] Radu Gramatovici. An efficient parser for a class of contextual languages. Fundamenta Informaticae, 33(3):211–238, March 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ilie:1998:GFW

[395] Lucian Ilie. Generalized factors of words. Fundamenta Informaticae, 33 (3):239–247, March 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kudlek:1998:ALR

[396] Manfred Kudlek and Alexandru Mateescu. Algebraic, linear and rational languages defined by mix operation. Fundamenta Informaticae, 33(3):249– 264, March 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Martinek:1998:LPG

[397] Pavel Martinek. Limits of pure grammars with monotone productions. Fundamenta Informaticae, 33(3):265–280, March 1998. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mitrana:1998:PCG

[398] Victor Mitrana. Parallelism in contextual grammars. Fundamenta Informaticae, 33(3):281–294, March 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Novotny:1998:RP

[399] Miroslav Novotný. Reduction of pregrammars. Fundamenta Informaticae, 33(3):295–307, March 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hindley:1998:P

[400] Roger Hindley. Preface. Fundamenta Informaticae, 33(4):??, December 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Brandt:1998:CAR

[401] Michael Brandt and Fritz Henglein. Coinductive axiomatization of recursive type equality and subtyping. Fundamenta Informaticae, 33(4): 309–338, December 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Brauner:1998:SAC

[402] Torben Braüner. A simple adequate categorical model for PCF, II. Fundamenta Informaticae, 33(4):339–368, December 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stark:1998:NER

[403] Ian Stark. Names, equations, relations: Practical ways to reason about new. Fundamenta Informaticae, 33(4):369–396, December 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Takeuti:1998:ASP

[404] Izumi Takeuti. An axiomatic system of parametricity. Fundamenta Informaticae, 33(4):397–432, December 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ferrari:1998:PSO

[405] Gian Luigi Ferrari and Ugo Montanari. Parameterized structured operational semantics. Fundamenta Informaticae, 34(1–2):1–31, May 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gomolinska:1998:CIM

[406] Anna Gomolińska. Credibility of information for modelling belief state and its change. Fundamenta Informaticae, 34(1–2):33–51, May 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

vanLinder:1998:FAO

[407] B. van Linder, W. van der Hoek, and J.-J. Ch. Meyer. Formalising abilities and opportunities of agents. Fundamenta Informaticae, 34(1–2):53–101, May 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

LaTorre:1998:SLI

[408] Salvatore La Torre, Margherita Napoli, and Domenico Parente. Synchronization of a line of identical processors at a given time. Fundamenta Informaticae, 34(1–2):103–128, May 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nguyen:1998:PED

[409] Sinh Hoa Nguyen and Hung Son Nguyen. Pattern extraction from data. Fundamenta Informaticae, 34(1–2):129–144, May 1998. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Son:1998:OHO

[410] Nguyen Hung Son. From optimal hyperplanes to optimal decision trees. Fundamenta Informaticae, 34(1–2):145–174, May 1998. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vagvolgyi:1998:GTT

[411] Sandor Vágvölgyi. The ground tree transducer game. Fundamenta Informaticae, 34(1–2):175–201, May 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Apt:1998:GEP

[412] Krzysztof R. Apt and Claude Kirchner. Guest Editors’ preface — foundations of constraint programming. Fundamenta Informaticae, 34(3):??, June 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kowalski:1998:ESL

[413] Robert A. Kowalski, Francesca Toni, and Gerhard Wetzel. Executing suspended logic programs. Fundamenta Informaticae, 34(3):203–224, June 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kirchner:1998:RBC

[414] Claude Kirchner and Christophe Ringeissen. Rule-based constraint programming. Fundamenta Informaticae, 34(3):225–262, June 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Castro:1998:BCS

[415] Carlos Castro. Building constraint satisfaction problem solvers using rewrite rules and strategies. Fundamenta Informaticae, 34(3):263–293, June 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Apt:1998:PTV

[416] Krzysztof R. Apt. A proof theoretic view of constraint programming. Fundamenta Informaticae, 34(3):295–321, June 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rety:1998:DCC

[417] Jean-Hugues Réty. Distributed concurrent constraint programming. Fundamenta Informaticae, 34(3):323–346, June 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skarbek:1998:PSI

[418] Władysław Skarbek. Preface: Special issue on image compression. Fundamenta Informaticae, 34(4):??, July 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cieplinski:1998:RIV

[419] Leszek Cieplinski. A review of image and video coding standards. Fundamenta Informaticae, 34(4):347–367, July 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Przelaskowski:1998:DPW

[420] Artur Przelaskowski. Detail preserving wavelet-based compression with adaptive context-based quantisation. Fundamenta Informaticae, 34(4): 369–388, July 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cieplinski:1998:ECM

[421] Leszek Cieplinski. Entropy-constrained multiresolution vector quantisation for image coding. Fundamenta Informaticae, 34(4):389–396, July 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rak:1998:VQ

[422] Remigiusz J. Rak. Vector quantisation. Fundamenta Informaticae, 34(4): 397–412, July 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mitra:1998:SPI

[423] Suman K. Mitra, C. A. Murthy, and Malay K. Kundu. A study on partitioned iterative function systems for image compression. Fundamenta Informaticae, 34(4):413–428, July 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skarbek:1998:AFO

[424] Władysław Skarbek. Analysis of fractal operator convergence by graph methods. Fundamenta Informaticae, 34(4):429–440, July 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sikora:1998:SAO

[425] Radosław Sikora and Władysław Skarbek. Stability analysis of Oja–RLS learning rule. Fundamenta Informaticae, 34(4):441–453, July 1998. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rakowski:1998:DFW

[426] Waldemar Rakowski and Zbigniew Bartosiewicz. Daubechies filters in wavelet image compression. Fundamenta Informaticae, 34(4):455–467, July 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wojdyllo:1998:WMM

[427] Piotr Wojdyłło. Wavelets and Mallat’s multiresolution analysis. Fundamenta Informaticae, 34(4):469–474, July 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Eiben:1998:P

[428] A. E. Eiben and Zbigniew Michalewicz. Preface. Fundamenta Informaticae, 35(1–4):??, January 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fogel:1998:UFH

[429] David B. Fogel. Unearthing a fossil from the history of evolutionary computation. Fundamenta Informaticae, 35(1–4):1–16, January 1998. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Esquivel:1998:SME

[430] Susana C. Esquivel, Héctor Ariel Leiva, and Raúl H. Gallardt. Selection mechanisms in evolutionary algorithms. Fundamenta Informaticae, 35(1– 4):17–33, January 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Eiben:1998:EEE

[431] A. E. Eiben and C. A. Schippers. On evolutionary exploration and exploitation. Fundamenta Informaticae, 35(1–4):35–50, January 1998. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Back:1998:OPC

[432] Thomas Bäck. An overview of parameter control methods by selfadaptation in evolutionary algorithms. Fundamenta Informaticae, 35(1– 4):51–66, January 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rudolph:1998:FMC

[433] Günter Rudolph. Finite Markov chain results in evolutionary computation: A tour d’horizon. Fundamenta Informaticae, 35(1–4):67–89, January 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Murthy:1998:OST

[434] C. A. Murthy, Dinabandhu Bhandari, and Sankar K. Pal. ϵ-optimal stopping time for genetic algorithms. Fundamenta Informaticae, 35(1–4):91– 111, January 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Khouja:1998:EAE

[435] Moutaz Khouja, Zbigniew Michalewicz, and Poorani Vijayaragavan. Evolutionary algorithm for economic lot and delivery scheduling problem. Fundamenta Informaticae, 35(1–4):113–123, January 1998. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sebag:1998:RME

[436] Michèle Sebag, Marc Schoenauer, and Mathieu Peyral. Revisiting the memory of evolution. Fundamenta Informaticae, 35(1–4):125–162, January 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Giordana:1998:LCP

[437] A. Giordana and G. Lo Bello. Learning classification programs: The genetic algorithm approach. Fundamenta Informaticae, 35(1–4):163–177, January 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Angeline:1998:HPE

[438] Peter J. Angeline. A historical perspective on the evolution of executable structures. Fundamenta Informaticae, 35(1–4):179–195, January 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Boryczka:1998:SAA

[439] Mariusz Boryczka. Some aspects of ant systems for the TSP. Fundamenta Informaticae, 35(1–4):197–209, January 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Seredynski:1998:NTP

[440] Franciszek Seredyński. New trends in parallel and distributed evolutionary computing. Fundamenta Informaticae, 35(1–4):211–230, January 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Deaton:1998:DCR

[441] Russell Deaton, Max Garzon, John Rose, D. R. Franceschetti, and S. E. Stevens. DNA computing: A review. Fundamenta Informaticae, 35(1– 4):231–245, January 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

DeJong:1998:ECW

[442] Kenneth De Jong. Evolutionary computation: Where we are and where we’re headed. Fundamenta Informaticae, 35(1–4):247–259, January 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bhat:1998:APO

[443] Girish Bhat and Doron Peled. Adding partial orders to linear temporal logic. Fundamenta Informaticae, 36(1):1–21, September 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Doherty:1998:GDC

[444] Patrick Doherty, Witold Łukaszewicz, and Andrzej Szałas. General domain circumscription and its effective reductions. Fundamenta Informaticae, 36(1):23–55, September 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rahonis:1998:SSS

[445] George Rahonis and Kai Salomaa. On the size of stack and synchronization alphabets of tree automata. Fundamenta Informaticae, 36(1):57–69, September 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Winkowski:1998:PCN

[446] Józef Winkowski. Processes of contextual nets and their characteristics. Fundamenta Informaticae, 36(1):71–101, September 1998. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Amin:1998:RRB

[447] Mohamed Amin. Remarks on the relationship between cooperating distributed grammar systems and Lindenmayer systems. Fundamenta Informaticae, 36(2–3):103–107, November 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Analyti:1998:SSN

[448] Anastasia Analyti, Nicolas Spyratos, and Panos Constantopoulos. On the semantics of a semantic network. Fundamenta Informaticae, 36(2–3): 109–144, November 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Berard:1998:CEP

[449] Béatrice Bérard, Antoine Petit, Volker Diekert, and Paul Gastin. Characterization of the expressive power of silent transitions in timed automata. Fundamenta Informaticae, 36(2–3):145–182, November 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dassow:1998:ROS

[450] Jürgen Dassow and Gheorghe Păun. Remarks on operations suggested by mutations in genomes. Fundamenta Informaticae, 36(2–3):183–200, November 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dovier:1998:UAV

[451] Agostino Dovier, Alberto Policriti, and Gianfranco Rossi. A uniform axiomatic view of lists, multisets, and sets, and the relevant unification algorithms. Fundamenta Informaticae, 36(2–3):201–234, November 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Middelburg:1998:TDC

[452] Kees A. Middelburg. Truth of duration calculus formulae in timed frames. Fundamenta Informaticae, 36(2–3):235–263, November 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stepaniuk:1998:ORS

[453] Jaroslaw Stepaniuk. Optimizations of rough set model. Fundamenta Informaticae, 36(2–3):265–283, November 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhiltsova:1998:EOC

[454] Larissa Zhiltsova. On entropy and optimal coding cost for stochastic language. Fundamenta Informaticae, 36(2–3):285–305, November 1998. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Burago:1998:ACC

[455] Dmitri Burago and Michel de Rougemont. On the average-case complexity of the graph reliability problem on Gaussian distributions. Fundamenta Informaticae, 36(4):307–315, December 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czech:1998:RPS

[456] Zbigniew J. Czech and Wojciech Mikanik. Randomized PRAM simulation. Fundamenta Informaticae, 36(4):317–336, December 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ostravsky:1998:PVS

[457] Jan Ostravský. Pregrammars versus submonoids. Fundamenta Informaticae, 36(4):337–343, December 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skarbek:1998:MOR

[458] Władysław Skarbek, Adam Pietrowcew, and Radosław Sikora. The modified oja-RLS algorithm, stochastic convergence analysis and application for image compression. Fundamenta Informaticae, 36(4):345–365, December 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vagvolgyi:1998:SMG

[459] Sándor Vágvölgyi. Simple and minimal ground term equation systems. Fundamenta Informaticae, 36(4):367–413, December 1998. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pal:1999:PSI

[460] Sankar K. Pal. Preface: Special issue on soft computing. Fundamenta Informaticae, 37(1–2):??, January 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dubois:1999:UPL

[461] Didier Dubois, Daniel Le Berre, Henri Prade, and Régis Sabbadin. Using possibilistic logic for modeling qualitative decision: ATMS-based algorithms. Fundamenta Informaticae, 37(1–2):1–30, January 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sinha:1999:CDI

[462] Sitabhra Sinha. Chaotic dynamics in iterated map neural networks with piecewise linear activation function. Fundamenta Informaticae, 37(1–2): 31–50, January 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kubota:1999:EMV

[463] Naoyuki Kubota and Toshio Fukuda. Ecological model of virusevolutionary genetic algorithm. Fundamenta Informaticae, 37(1–2):51– 70, January 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tanaka:1999:IRA

[464] Hideo Tanaka and Haekwan Lee. Interval regression analysis with polynomials and its similarity to rough sets concept. Fundamenta Informaticae, 37(1–2):71–87, January 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Friedman:1999:FIE

[465] Menahem Friedman, Ma Ming, and Abraham Kandel. On fuzzy integral equations. Fundamenta Informaticae, 37(1–2):89–99, January 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

DiGesu:1999:AVS

[466] Vito Di Gesù. Artificial vision and soft computing. Fundamenta Informaticae, 37(1–2):101–119, January 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sastry:1999:DEB

[467] K. K. N. Sastry, L. Behera, and I. J. Nagrath. Differential evolution based fuzzy logic controller for nonlinear process control. Fundamenta Informaticae, 37(1–2):121–136, January 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Karayiannis:1999:RLV

[468] Nicolaos B. Karayiannis. Reformulating learning vector quantization and radial basis neural networks. Fundamenta Informaticae, 37(1–2):137–175, January 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bandyopadhyay:1999:RBV

[469] Sanghamitra Bandyopadhyay and Sankar K. Pal. Relation between VGAclassifier and MLP: Determination of network architecture. Fundamenta Informaticae, 37(1–2):177–199, January 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Burkhard:1999:P

[470] Hans-Dieter Burkhard, Peter Starke, and Ludwik Czaja. Preface. Fundamenta Informaticae, 37(3):??, February 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Baar:1999:IDV

[471] Thomas Baar and Ekkart Kindler. ILF and DAWN for verifying distributed algorithms — an idea for a tool. Fundamenta Informaticae, 37 (3):201–211, February 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:1999:NDP

[472] Ludwik Czaja. Net-definability of process languages. Fundamenta Informaticae, 37(3):213–223, February 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Farwer:1999:LLV

[473] Berndt Farwer. A linear logic view of object Petri nets. Fundamenta Informaticae, 37(3):225–246, February 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Heljanko:1999:ULP

[474] Keijo Heljanko. Using logic programs with stable model semantics to solve deadlock and reachability problems for 1-safe Petri nets. Fundamenta Informaticae, 37(3):247–268, February 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Junttila:1999:FSA

[475] Tommi A. Junttila. Finding symmetries of algebraic system nets. Fundamenta Informaticae, 37(3):269–289, February 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kindler:1999:IDA

[476] Ekkart Kindler and Sibylle Peuker. Integrating distributed algorithms into distributed systems. Fundamenta Informaticae, 37(3):291–309, February 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Popova-Zeugmann:1999:APT

[477] Louchka Popova-Zeugmann and Dirk Schlatter. Analyzing paths in time Petri nets. Fundamenta Informaticae, 37(3):311–327, February 1999. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bohm:1999:FPE

[478] Corrado Böhm. Fixed point equations inside the algebra of normal forms. Fundamenta Informaticae, 37(4):329–342, March 1999. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:1999:RHS

[479] Ludwik Czaja. Representing hand-shake channel communication in the calculus of cause-effect structures. Fundamenta Informaticae, 37(4):343– 368, March 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jung:1999:MLS

[480] Achim Jung, Mathias Kegelmann, and M. Andrew Moshier. Multi lingual sequent calculus and coherent spaces. Fundamenta Informaticae, 37(4): 369–412, March 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kudlek:1999:ILC

[481] Manfred Kudlek. Iteration lemmata for certain classes of word, trace and graph languages. Fundamenta Informaticae, 37(4):413–422, March 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Karhumaki:1999:P

[482] J. Karhumäki, A. Mateescu, and G. Rozenberg. Preface. Fundamenta Informaticae, 38(1–2):??, April 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ben-Amram:1999:PVT

[483] Amir M. Ben-Amram and Neil D. Jones. A precise version of a time hierarchy theorem. Fundamenta Informaticae, 38(1–2):1–15, April 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Penttonen:1999:RCR

[484] Martti Penttonen. Relative cost random access machines. Fundamenta Informaticae, 38(1–2):17–23, April 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skyum:1999:CIL

[485] Sven Skyum, Gudmund S. Frandsen, Peter Bro Miltersen, and Peter G. Binderup. The complexity of identifying large equivalence classes. Fundamenta Informaticae, 38(1–2):25–37, April 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wanne:1999:GMA

[486] Merja Wanne and Matti Linna. A general model for adjacency. Fundamenta Informaticae, 38(1–2):39–50, April 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bogdanovic:1999:TCE

[487] S. Bogdanović, M. Ćirić, T. Petković, B. Imreh, and M. Steinby. Traps, cores, extensions and subdirect decompositions of unary algebras. Fundamenta Informaticae, 38(1–2):51–60, April 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nielsen:1999:BOM

[488] Mogens Nielsen and Thomas Hune. Bisimulation and open maps for timed transition systems. Fundamenta Informaticae, 38(1–2):61–77, April 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jurvanen:1999:DSU

[489] Eija Jurvanen and Marjo Lipponen. Distinguishability, simulation and universality of Moore tree automata. Fundamenta Informaticae, 38(1–2): 79–91, April 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kari:1999:CDS

[490] Jarkko Kari. On the circuit depth of structurally reversible cellular automata. Fundamenta Informaticae, 38(1–2):93–107, April 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ilie:1999:SPF

[491] Lucian Ilie. Subwords and power-free words are not expressible by word equations. Fundamenta Informaticae, 38(1–2):109–118, April 1999. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Karhumaki:1999:DTT

[492] Juhani Karhumäki and Sabrina Mantaci. Defect theorems for trees. Fundamenta Informaticae, 38(1–2):119–133, April 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ruohonen:1999:SER

[493] Keijo Ruohonen. Solving equivalence of recurrent sequences in groups by polynomial manipulation. Fundamenta Informaticae, 38(1–2):135–148, April 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ding:1999:CEA

[494] Cunsheng Ding, Kwok Yan Lam, and Chaoping Xing. Construction and enumeration of all binary duadic codes of length p^{m}. Fundamenta Informaticae, 38(1–2):149–161, April 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Koskinen:1999:NIK

[495] Jukka A. Koskinen. Non-injectivity and knapsacks. Fundamenta Informaticae, 38(1–2):163–180, April 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Niemi:1999:SZK

[496] Valtteri Niemi and Ari Renvall. Solitaire zero-knowledge. Fundamenta Informaticae, 38(1–2):181–188, April 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Halava:1999:UIW

[497] Vesa Halava and Tero Harju. Undecidability in integer weighted finite automata. Fundamenta Informaticae, 38(1–2):189–200, April 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Honkala:1999:EPD

[498] Juha Honkala. The equivalence problem of D0L and DF0L power series. Fundamenta Informaticae, 38(1–2):201–208, April 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Turakainen:1999:DSP

[499] Paavo Turakainen. On the decidability of some problems concerning morphisms, regular sets, and post correspondence problem. Fundamenta Informaticae, 38(1–2):209–221, April 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Janicki:1999:CSN

[500] Ryszard Janicki and Maciej Koutny. On causality semantics of nets with priorities. Fundamenta Informaticae, 38(3):223–255, May 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kappes:1999:MBC

[501] Martin Kappes. Multi-bracketed contextual rewriting grammars. Fundamenta Informaticae, 38(3):257–280, May 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Koutras:1999:MVM

[502] Costas D. Koutras, George Koletsos, and Stathis Zachos. Many-valued modal non-monotonic reasoning: Sequential stable sets and logics with linear truth spaces. Fundamenta Informaticae, 38(3):281–324, May 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bono:1999:SEI

[503] Viviana Bono, Michele Bugliesi, Mariangiola Dezani-Ciancaglini, and Luigi Liquori. A subtyping for extensible, incomplete objects. Fundamenta Informaticae, 38(4):325–364, December 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bordihn:1999:APL

[504] Henning Bordihn, Henning Fernau, and Markus Holzer. On accepting pure Lindenmayer systems. Fundamenta Informaticae, 38(4):365–375, December 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Corradini:1999:RAF

[505] Flavio Corradini, Domenicantonio D’Ortenzio, and Paola Inverardi. On the relationships among four timed process algebras. Fundamenta Informaticae, 38(4):377–395, December 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paun:1999:SS

[506] Gheorghe Păun and Sheng Yu. On synchronization in P systems. Fundamenta Informaticae, 38(4):397–410, December 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Plaza:1999:P

[507] Jan A. Plaza and John A. Campbell. Preface. Fundamenta Informaticae, 39(1–2):??, July 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ballarin:1999:PAE

[508] Clemens Ballarin and Lawrence C. Paulson. A pragmatic approach to extending provers by computer algebra — with applications to coding theory. Fundamenta Informaticae, 39(1–2):1–20, July 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Benhamou:1999:HMF

[509] Belaid Benhamou and Laurent Henocque. A hybrid method for finite model search in equational theories. Fundamenta Informaticae, 39(1–2): 21–38, July 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bertoli:1999:SIT

[510] P. G. Bertoli, J. Calmet, F. Giunchiglia, and K. Homann. Specification and integration of theorem provers and computer algebra systems. Fundamenta Informaticae, 39(1–2):39–57, July 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Egly:1999:IPT

[511] Uwe Egly and Stephan Schmitt. On intuitionistic proof transformations, their complexity, and application to constructive program synthesis. Fundamenta Informaticae, 39(1–2):59–83, July 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fevre:1999:CCA

[512] Stéphane Fèvre and Dongming Wang. Combining Clifford algebraic computing and term-rewriting for geometric theorem proving. Fundamenta Informaticae, 39(1–2):85–104, July 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fitting:1999:QML

[513] Melvin Fitting. On quantified modal logic. Fundamenta Informaticae, 39 (1–2):105–121, July 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fuchs:1999:USC

[514] Dirk Fuchs. On the use of subgoal clauses in bottom-up and top-down calculi. Fundamenta Informaticae, 39(1–2):123–143, July 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lynch:1999:BCC

[515] Christopher Lynch and Christelle Scharff. Basic completion with E-cycle simplification. Fundamenta Informaticae, 39(1–2):145–165, July 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Monfroy:1999:OAF

[516] Eric Monfroy and Christophe Ringeissen. An open automated framework for constraint solver extension: the SoleX approach. Fundamenta Informaticae, 39(1–2):167–187, July 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pientka:1999:AIS

[517] Brigitte Pientka and Christoph Kreitz. Automating inductive specification proofs. Fundamenta Informaticae, 39(1–2):189–209, July 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sendra:1999:ARR

[518] J. Rafael Sendra and Franz Winkler. Algorithms for rational real algebraic curves. Fundamenta Informaticae, 39(1–2):211–228, July 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Duntsch:1999:RAQ

[519] Ivo Düntsch, Hui Wang, and Stephen McCloskey. Relations algebras in qualitative spatial reasoning. Fundamenta Informaticae, 39(3):229–248, August 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dudakov:1999:CPM

[520] Sergey M. Dudakov. On the complexity of perfect models of logic programs. Fundamenta Informaticae, 39(3):249–258, August 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Heinemann:1999:CCM

[521] Bernhard Heinemann. The complexity of certain modal formulas on binary ramified subset trees. Fundamenta Informaticae, 39(3):259–272, August 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Intrigila:1999:CSM

[522] B. Intrigila, P. Inverardi, and M. Venturini Zilli. A comprehensive setting for matching and unification over iterative terms. Fundamenta Informaticae, 39(3):273–304, August 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Knapik:1999:TST

[523] Teodor Knapik and Hugues Calbrix. Thue specifications and their monadic second-order properties. Fundamenta Informaticae, 39(3):305– 325, August 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Muller:1999:AA

[524] Heike Müller, Torsten Hilbrich, and Ralf Kühnel. An assistant agent. Fundamenta Informaticae, 39(3):327–336, August 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chikalov:1999:ATC

[525] Igor Chikalov. On average time complexity of decision trees and branching programs. Fundamenta Informaticae, 39(4):337–357, September 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ejdys:1999:MWL

[526] Piotr Ejdys and Grzegorz Góra. The more we learn the less we know? Fundamenta Informaticae, 39(4):359–374, September 1999. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hage:1999:MPS

[527] Jurriaan Hage. The membership problem for switching classes with skew gains. Fundamenta Informaticae, 39(4):375–387, September 1999. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Marek:1999:CTR

[528] V. Wiktor Marek and Mirosław Truszczyński. Contributions to the theory of rough sets. Fundamenta Informaticae, 39(4):389–409, September 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Woltert:1999:MDL

[529] Prank Woltert and Michael Zakharyaschev. Modal description logics: modalizing roles. Fundamenta Informaticae, 39(4):411–438, September 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ambainis:1999:IIP

[530] Andris Ambainis, Rūsiņš Freivalds, and Carl H. Smith. Inductive inference with procrastination: Back to definitions. Fundamenta Informaticae, 40 (1):1–16, October 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Doherty:1999:MQD

[531] Patrick Doherty, Jarosław Kachniarz, and Andrzej Szałas. Meta-queries on deductive databases. Fundamenta Informaticae, 40(1):17–30, October 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Johnson:1999:PDD

[532] C. A. Johnson. Processing deductive databases under the disjunctive stable model semantics. Fundamenta Informaticae, 40(1):31–51, October 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Preoteasa:1999:RBU

[533] Viorel Preoteasa. A relation between unambiguous regular expressions and abstract data types. Fundamenta Informaticae, 40(1):53–77, October 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Su:1999:CES

[534] Kaile Su and Wei Li. Computation of extensions of seminormal default theories. Fundamenta Informaticae, 40(1):79–102, October 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Janicki:1999:P

[535] Ryszard Janicki. Preface. Fundamenta Informaticae, 40(2–3):??, November 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Koutny:1999:PLC

[536] Maciej Koutny. Peter Lauer and COSY. Fundamenta Informaticae, 40(2– 3):103–107, November 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Arnold:1999:AFD

[537] André Arnold, Gérald Point, Alain Griffault, and Antoine Rauzy. The AltaRica formalism for describing concurrent systems. Fundamenta Informaticae, 40(2–3):109–124, November 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Best:1999:GCO

[538] Eike Best and Alexander Lavrov. Generalised composition operations for high-level Petri nets. Fundamenta Informaticae, 40(2–3):125–163, November 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Busi:1999:PSP

[539] Nadia Busi and G. Michele Pinna. Process semantics for place/transition nets with inhibitor and read arcs. Fundamenta Informaticae, 40(2–3): 165–197, November 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Devillers:1999:GRT

[540] Raymond Devillers and Joël Goossens. General response time computation for the deadline driven scheduling of periodic tasks. Fundamenta Informaticae, 40(2–3):199–219, November 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maggiolo-Schettini:1999:ATA

[541] Andrea Maggiolo-Schettini and Simone Tini. Applying techniques of asynchronous concurrency to synchronous languages. Fundamenta Informaticae, 40(2–3):221–250, November 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pietkiewicz-Koutny:1999:SPE

[542] Marta Pietkiewicz-Koutny. The synthesis problem for elementary net systems with inhibitor arcs. Fundamenta Informaticae, 40(2–3):251–283, November 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Shields:1999:AVS

[543] M. W. Shields. An acceptance vector semantics for path programs. Fundamenta Informaticae, 40(2–3):285–316, November 1999. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

delCerro:1999:TBD

[544] Luis Fariñas del Cerro and Olivier Gasquet. Tableaux based decision procedures for modal logics of confluence and density. Fundamenta Informaticae, 40(4):317–333, December 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

denHartog:1999:FAM

[545] J. I. den Hartog, E. P. de Vink, and J. W. de Bakker. Full abstractness of a metric semantics for action refinement. Fundamenta Informaticae, 40 (4):335–382, December 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nguyen:1999:RSA

[546] Hung Son Nguyen and Sinh Hoa Nguyen. Rough sets and association rule generation. Fundamenta Informaticae, 40(4):383–405, December 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ordman:1999:GME

[547] E. T. Ordman, E. Eberbach, and A. Anwar. Generalized mutual exclusion with semaphores only. Fundamenta Informaticae, 40(4):407–422, December 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).